### v1

# Offline Clustering of Linear Bandits: Unlocking the Power of Clusters in Data-Limited Environments

Abstract
Contextual linear multi-armed bandits are a learning framework for making a sequence of decisions, e.g., advertising recommendations for a sequence of arriving users. Recent works have shown that clustering these users based on the similarity of their learned preferences can significantly accelerate the learning. However, prior work has primarily focused on the online setting, which requires continually collecting user data, ignoring the offline data widely available in many applications. To tackle these limitations, we study the offline clustering of bandits (Off-ClusBand) problem, which studies how to use the offline dataset to learn cluster properties and improve decision-making across multiple users. The key challenge in Off-ClusBand arises from data insufficiency for users: unlike the online case, in the offline case, we have a fixed, limited dataset to work from and thus must determine whether we have enough data to confidently cluster users together. To address this challenge, we propose two algorithms: Off-C2LUB, which we analytically show performs well for arbitrary amounts of user data, and Off-CLUB, which is prone to bias when data is limited but, given sufficient data, matches a theoretical lower bound that we derive for the offline clustered MAB problem. We experimentally validate these results on both real and synthetic datasets.
1 Introduction
Contextual bandits (Chu et al., 2011; Agrawal and Devanur, 2016) extend the traditional multi-armed bandit (MAB) framework (Robbins, 1952; Lai and Robbins, 1985; Auer et al., 2002) for optimizing online decision-making to settings where decision rewards depend on given contexts, often through a linear function. Recent works have proposed to accelerate the learning in contextual MAB by taking advantage of the clustered structure in many real-world applications, e.g., social networking (Buccapatnam et al., 2013; Delporte et al., 2013; Bitaghsir et al., 2019) and advertisement recommendation (Ek and Stigsson, 2015; Yan et al., 2022; Aramayo et al., 2023). This framework assumes that a given set of users arrives at the system in sequence; users generally arrive multiple times within the sequence (e.g., users visiting a website multiple times and being shown an ad at each visit). These users can be partitioned into clusters that share the same or similar preferences. The learning agent adaptively clusters users into groups based on their observed responses to the actions taken, leveraging the preference similarities within these groups to improve its decisions. This online and adaptive clustering approach enables more efficient and personalized decision-making, which has been extensively studied across diverse and practical settings (Gentile et al., 2014; Li et al., 2019; Ban et al., 2024). Detailed related works can be found in Appendix 2.
Despite the significant progress in clustering-based bandit frameworks, prior work has largely overlooked scenarios where offline data plays a crucial role in learning. For example, advertisement platforms often accumulate extensive historical data on the effectiveness of given ads across different users. This data can be used to pre-compute user clusters, enabling more informed and efficient future recommendations. MAB approaches have also been applied to choosing medical treatments (Komorowski et al., 2018; Zhou et al., 2019), for which patients with similar physiological profiles often exhibit comparable responses. However, ethical and practical constraints make it infeasible to directly experiment with treatments on patients. Instead, learning must be conducted using preexisting historical treatment data. To address this gap, we introduce, to the best of our knowledge, the first algorithms for offline clustered bandits. Our approach leverages offline data to identify clusters and optimize decision-making in settings where online exploration is constrained or infeasible, paving the way for broader applications of MAB-based decision-making.
Extending bandit clustering algorithms from the online to the offline setting poses a fundamentally new challenge: unlike in the online setting, where data continually arrives during algorithm execution, the offline setting relies on a fixed offline training dataset. The continual arrival of new data in the online setting allows the algorithm to iteratively refine its understanding of the clustering structure; indeed, over $T$ rounds, only $O(\log(T))$ rounds are typically needed to accurately identify clusters. However, in the offline setting, the number of samples per user is inherently limited. For instance, some users may have sparse historical data due to infrequent activity, or data on certain treatments or actions may be scarce. This limitation creates significant challenges in accurately estimating the preferences of users with limited data and thus in clustering users with similar preferences. We must thus develop criteria to confidently identify similar users, and then account for potential clustering errors—caused by limited or noisy data—in designing and analyzing the bandit algorithm. Unlike standard offline statistical learning (Zhang and Zhang, 2014; Cai and Guo, 2017), where the focus is on estimating parameters with confidence, our setting also requires making structural decisions—such as whether to group users together—based on different amounts of data as well as confidence levels, which fundamentally complicates both the algorithm and its analysis.
To address this challenge, we introduce the Offline Clustering of Bandits (Off-ClusBand) problem. In Off-ClusBand, there are $U$ users represented by the set $\mathcal{U}$, each associated with a preference vector. The users are partitioned into $J$ clusters, with users in the same cluster sharing an identical preference vector. An input dataset $\mathcal{D}$ is provided, containing historical samples of the actions taken and rewards received, assumed to be a linear function of user actions and preference vectors. The objective is to design an algorithm that minimizes the suboptimality gap—the difference between the reward achieved by the algorithm and the optimal reward achievable in hindsight—for any given test user $u_{\text{test}}\in\mathcal{U}$ and an arbitrary action set $A_{\text{test}}$, which may not have been observed in the offline dataset.
Contributions. (i) In Section 3, we formalize the Off-ClusBand problem. We then design the Off-C2LUB algorithm (Algorithm 1) in Section 4, which addresses the challenge of inaccurate clustering due to limited user data by introducing a parameter $\hat{\gamma}$ that thresholds similar users. In Theorem 4.2, we upper bound Off-C2LUB’s suboptimality gap, accounting for two types of error. The homogeneous error decreases as more users sharing the same preference vector of the test user are identified, and this reduction accelerates with an increase in $\hat{\gamma}$ (i.e., the threshold for declaring users to be similar). Conversely, the heterogeneous error, which comes from erroneously treating users with distinct preferences as similar to the test user, may increase with $\hat{\gamma}$. By selecting $\hat{\gamma}\leq\gamma$, where $\gamma$ is the minimum gap between any two users from different clusters, results in zero heterogeneous error but higher homogeneous error; while selecting a large $\hat{\gamma}$ can cause a substantial increase in heterogeneous error but evidently decrease the homogeneous error. We analyze the influence of $\hat{\gamma}$ and provide policies for selecting it under different scenarios, nearly matching the lower bound of the suboptimality gap that we derive in Section 6.
(ii) We then present the Off-CLUB algorithm (Algorithm 2) in Section 5, which does not require $\hat{\gamma}$. Assuming sufficient user data for accurate clustering, we upper bound Off-CLUB’s suboptimality in Theorem 5.2. However, Off-CLUB struggles when the dataset is sparse or lacks sufficient samples. The recommended algorithms based on different data conditions is provided in Table 1.
(iii) Finally, in Section 7, we validate our algorithms through both synthetic and real-world experiments. These experiments demonstrate the strong performance for the suboptimality gap of Off-C2LUB across both scenarios with different amounts of data, especially when coupled with our proposed $\hat{\gamma}$-selection policies. Off-C2LUB reduces the suboptimality gap by 50% compared to previous offline clustering algorithms. Moreover, they confirm the effectiveness of Off-CLUB under sufficient data and highlight its limitations when data is insufficient.
2 Related Works
Table 1: Recommended Algorithms Based on Data Conditions
Dataset $\gamma$ Known Unknown
Sufficient Equivalent Off-CLUB
Insufficient Off-C2LUB Off-C2LUB
Offline Bandits. Offline multi-armed bandits (MAB), introduced by  Shivaswamy and Joachims (2012), are a specialized variant of offline reinforcement learning (RL) that utilize pre-collected data for decision-making based on confidence balls for regression problems (Javanmard and Montanari, 2014; Zhang and Zhang, 2014; Cai and Guo, 2017). These methods have been applied in domains such as robotics  (Kumar et al., 2020), autonomous driving  (Yurtsever et al., 2020), and education  (Singla et al., 2021). Some offline MAB studies consider scenarios where the offline distribution remains consistent with the online rewards  (Bu et al., 2020; Banerjee et al., 2022), while others address settings with reward distribution shifts  (Zhang et al., 2019; Cheung and Lyu, 2024). Research has also explored offline data provided sequentially  (Gur and Momeni, 2022) or in distinct groups  (Bouneffouf et al., 2019; Ye et al., 2020; Tennenholtz et al., 2021). Among these,  Li et al. (2022); Wang et al. (2024) focus on offline contextual linear bandits, which are most closely aligned with our work. Recent works have explored offline RL with homogeneous agents  (Woo et al., 2023, 2024), as well as settings involving heterogeneous agents  (Zhou et al., 2024), where each agent serves a heterogeneous user, or heterogeneous data resources  (Wang et al., ; Shi et al., 2023). Our work addresses offline bandits with multiple data resources for heterogeneous users and introduces a technique to identify and merge similar datasets for more accurate predictions, setting it apart from previous studies.
Online Clustering of Bandits. Online clustering of bandits have been extensively studied across various settings and applications, beginning with Gentile et al. (2014) and including studies linking clusters to items for recommendations (Li et al., 2016a, b; Gentile et al., 2017) and exploring non-linear reward functions (Nguyen et al., 2020; Ban et al., 2024). Distributed bandit clustering has been a focus in  (Korda et al., 2016; Cherkaoui et al., 2023), while extensions to dependent arms were explored in  (Wang et al., 2019). Several works focus on the way to simplify the basic assumptions (Li et al., 2019, 2025). Applications range from recommendation systems  (Yang and Toni, 2018; Yan et al., 2022) to vehicular edge computing and dynamic pricing  (Lin et al., 2022; Miao et al., 2022). Recent studies have also addressed privacy concerns  (Liu et al., 2022), dynamic cluster switching  (Nguyen and Lauw, 2014), unified clustering with non-stationary bandits  (Li et al., 2021), corrupted user detection (Wang et al., 2023) and conversational  (Wu et al., 2021; Li et al., 2023; Dai et al., 2024a, b). In contrast to the online setting, where infinite samples can be drawn from unlimited online data, a key challenge in our setting lies in addressing scenarios with limited offline data. To the best of our knowledge, this is the first work to explore clustering bandits specifically in the offline bandit learning setting.
3 Problem Formulation
In this section, we introduce the setting for the “Offline Clustering of Bandits (Off-ClusBand)” problem. We consider $U$ users, represented as $\mathcal{U}=\{1,\cdots,U\}$, where each user $u$ is associated with a preference vector $\bm{\theta}_{u}$. To capture the heterogeneity within the user population, users are grouped into $J(\leq U)$ disjoint clusters, with all users in the same cluster sharing an identical preference vector. Precisely, we define $\mathcal{V}=\bigcup_{j=1}^{J}\mathcal{V}(j)$, where $\mathcal{V}(j)$ denotes the set of users within cluster $j$, hence $\mathcal{V}(j)\cap\mathcal{V}(j^{\prime})=\emptyset$ for $j\neq j^{\prime}$. Additionally, $\bm{\theta}_{u}=\bm{\theta}_{u^{\prime}}$ if and only if there exists a $j\in[J]$ such that $\{u,u^{\prime}\}\subseteq\mathcal{V}(j)$. The cluster to which user $u$ belongs is denoted by $j_{u}$, and the shared preference vector for users in cluster $\mathcal{V}(j)$ is denoted as $\bm{\theta}^{j}$. Notably, both the actual partition of clusters and the number of clusters remain unknown to the users. For user $u$, we refer to users in the same cluster as homogeneous users, and those in different clusters as heterogeneous users.
In the offline setting, we are given a dataset $\mathcal{D}=\{\mathcal{D}_{1},\cdots,\mathcal{D}_{U}\}$ comprising offline samples collected from $U$ users. For each user $u$, we observe $N_{u}$ samples, forming $\mathcal{D}_{u}=\{(\bm{a}_{u}^{i},r_{u}^{i})\}_{i=1}^{N_{u}}$, where $\{\bm{a}_{u}^{i}\}_{i=1}^{N_{u}}$ represents the $N_{u}$ given actions, and $\{r_{u}^{i}\}_{i=1}^{N_{u}}$ denotes the corresponding rewards, independently drawn as $r_{u}^{i}\sim R(\bm{a}_{u}^{i})$. Here, $R(\bm{a}_{u}^{i})$ is a 1-subgaussian reward distribution associated with the action $\bm{a}_{u}^{i}$, and $\mathbb{E}[r_{u}^{i}]=\langle\bm{\theta}_{u},\bm{a}_{u}^{i}\rangle$ represents the mean reward for user $u$ under action $\bm{a}_{u}^{i}$. Additionally, we assume that the actions adhere to the following assumption (Gentile et al., 2014, 2017; Blaser et al., 2024).
Assumption 3.1 (Action Regularity).
Each action $\bm{a}$ is independently and identically drawn from a fixed distribution $\rho$ over the set $\{\bm{a}\in\mathbb{R}^{d}:||\bm{a}||_{2}\leq 1\}$. The covariance matrix $\mathbb{E}_{\bm{a}\sim\rho}[\bm{a}\bm{a}^{\top}]$ is full rank with a minimum eigenvalue of $\lambda_{a}>0$.
Let $\pi$ denote an algorithm that selects an action from an action set $A_{\text{test}}\subseteq\mathbb{R}^{d}$ for any given user $u_{\text{test}}\in\mathcal{U}$. The goal in this setting is to design an algorithm $\pi$ that minimizes the suboptimality gap for an incoming user $u_{\text{test}}$ to be served, defined as (Jin et al., 2021; Li et al., 2022):
$\text{SubOpt}(u_{\text{test}},\pi,A_{\text{test}})\coloneqq\langle\bm{\theta}_{u_{\text{test}}},\bm{a}_{\text{test}}^{*}\rangle-\langle\bm{\theta}_{u_{\text{test}}},\bm{a}_{\pi(u_{\text{test}})}\rangle,$
where $\bm{a}_{\text{test}}^{*}$ and $\bm{a}_{\pi(u_{\text{test}})}$ represent the optimal (i.e., reward-maximizing) action in $A_{\text{test}}$ and the action selected by $\pi$ for $u_{\text{test}}$, respectively.
4 Algorithm for Generic: Offline Connection-based Clustering of Bandits
This section introduces our first algorithm: Offline Connection-based Clustering of Bandits Algorithm (Off-C2LUB) and a theoretical upper bound on its suboptimality.
4.1 Algorithm 1: Off-C2LUB
Compared to the online clustering of bandits problem, the core challenge in Off-ClusBand lies in data insufficiency. In the online setting, algorithms start with a complete graph over all users, where each node represents a user and each edge encodes a tentative assumption that the connected users are thought to be in the same cluster (Gentile et al., 2014; Li et al., 2019) and iteratively delete edges to separate users into clusters (i.e., remaining connected subgraphs). With abundant online data, this process accurately removes incorrect connecting edges. However, in the offline setting, limited data makes it difficult to precisely remove edges, often leading to the inclusion of heterogeneous users within the same cluster, which introduces bias and reduces decision quality. Therefore, building on the challenges described above, the core idea behind Off-C2LUB is to eliminate connections between users with significantly different preference vectors, i.e., prevent them from being assigned as neighbors. Specifically, the algorithm treats users with few samples independently, while leveraging the clustering property only for those with a sufficient dataset $\mathcal{D}_{u}$. The workflow of the proposed Off-C2LUB successfully avoids learning incorrect clusters by ensuring that predictions are made solely using data from users who share the same (or similar) preference vector as $u_{\textnormal{test}}$. The pseudo-code for Off-C2LUB, consisting a Cluster Phase and a Decision Phase, is presented in Algorithm 1.
Algorithm 1 Off-C2LUB
1:   Input: User $u_{\textnormal{test}}$, action set $A_{\textnormal{test}}$, dataset $\mathcal{D}$, parameters $\alpha>1$, $\lambda>0$, $\delta>0$ and $\hat{\gamma}\geq 0$
2:   Initialization: Construct a null graph $\mathcal{G}=(\mathcal{V},\emptyset)$, and compute $M_{u}$, $\bm{b}_{u}$, $\text{CI}_{u}$ as in Equation 1 and $\hat{\bm{\theta}}_{u}=M_{u}^{-1}\bm{b}_{u}$ for each user $u\in\mathcal{V}$
[…]
4:  for any two users $u_{1}$, $u_{2}\in\mathcal{V}$ do
5:      Connect $(u_{1},u_{2})$ if conditions in Equation 2 are satisfied
6:  end for
7:  Let $\mathcal{G}_{\hat{\gamma}}=(\mathcal{V},\mathcal{E}_{\hat{\gamma}})$ denote the graph afterwards
8:  for each user $u\in\mathcal{V}$ do
9:      Aggregate data and compute statistics:
$\textstyle\tilde{\mathcal{D}}_{u}=\bigcup_{(u,v)\in\mathcal{E}_{\hat{\gamma}}}\left\{(\bm{a}_{v}^{i},r_{v}^{i})\in\mathcal{D}_{v}\right\}\cup\mathcal{D}_{u},\ \tilde{n}_{u}=1+\sum_{v\in\mathcal{V}}\mathbb{I}\left[(u,v)\in\mathcal{E}_{\hat{\gamma}}\right]$
$\textstyle\tilde{M}_{u}=\lambda\tilde{n}_{u}\cdot I+\sum_{(\bm{a},r)\in\tilde{\mathcal{D}}_{u}}\bm{a}\bm{a}^{\top},\ \tilde{N}_{u}=\left|\tilde{\mathcal{D}}_{u}\right|,\ \tilde{\bm{b}}_{u}=\sum_{(\bm{a},r)\in\tilde{\mathcal{D}}_{u}}r\bm{a},\ \tilde{\bm{\theta}}_{u}=\left(\tilde{M}_{u}\right)^{-1}\tilde{\bm{b}}_{u}$
10:  end for
[…]
12:  Select $\bm{a}_{\textnormal{test}}=\mathop{\mathrm{argmax}}\limits_{\bm{a}\in A_{\textnormal{test}}}\left(\tilde{\bm{\theta}}_{u_{\textnormal{test}}}^{\top}\bm{a}-\beta\left\|\bm{a}\right\|_{\tilde{M}_{u_{\textnormal{test}}}^{-1}}\right)$ with $\beta=\sqrt{d\log\left(1+\frac{\tilde{N}_{u_{\textnormal{test}}}}{\lambda\tilde{n}_{u_{\textnormal{test}}}\cdot d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{\lambda}$
Input and Initialization. Algorithm 1 takes as input the test user $u_{\textnormal{test}}\in\mathcal{U}$, the action set $A_{\textnormal{test}}$, the dataset $\mathcal{D}$ and several parameters (Line 1), which will be discussed later. The algorithm initializes the following (Line 2) for each user $u$:
$M_{u}=\lambda I+\sum_{i=1}^{N_{u}}\bm{a}_{u}^{i}\left(\bm{a}_{u}^{i}\right)^{\top},\ \bm{b}_{u}=\sum_{i=1}^{N_{u}}r_{u}^{i}\bm{a}_{u}^{i},\,\text{CI}_{u}=\frac{\sqrt{d\log\left(1+\frac{N_{u}}{\lambda d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{\lambda}}{\sqrt{\lambda_{a}N_{u}/2}}.$ (1)
Here, $M_{u}$ represents Gramian matrix regularized by the input parameter $\lambda$, $\bm{b}_{u}$ is the moment matrix of the response variable, $\hat{\bm{\theta}}_{u}$ is the ridge regression estimate of $\bm{\theta}_{u}$ for user $u$, and $\text{CI}_{u}$ is the confidence interval bound. These computed values are critical for both phases of the algorithm.
Cluster Phase. Instead of starting with a complete graph as in the online setting, Algorithm 1 initializes a null graph ${\mathcal{G}}$ with only vertices but no edges. During the Cluster Phase, the algorithm iterates over all user pairs $u_{1}$ and $u_{2}$ in $\mathcal{V}$, connecting them if the following condition holds:
$\left\|\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}\right\|_{2}<\hat{\gamma}-\alpha\left(\text{CI}_{u_{1}}+\text{CI}_{u_{2}}\right)\text{ and }\min\left\{N_{u_{1}},N_{u_{2}}\right\}\geq\frac{16}{\lambda_{a}^{2}}\log\left(\frac{8Ud}{\lambda_{a}^{2}\delta}\right).$ (2)
The condition in Equation 2 ensures that the difference between the true preference vectors $\bm{\theta}_{u_{1}}$ and $\bm{\theta}_{u_{2}}$ is no greater than $\hat{\gamma}$ (see Theorem 4.2 for a detailed explanation). By only connecting users with similar estimated preference vectors, the algorithm prevents the introduction of bias that arises when users with highly divergent preference vectors are clustered together. Therefore, the choice of $\hat{\gamma}$ significantly affects the algorithm’s performance. As discussed in Section 4.2, a smaller $\hat{\gamma}$ helps avoid using samples from heterogeneous users of $u_{\text{test}}$, reducing bias, but may also limit access to samples from homogeneous users. In contrast, a larger $\hat{\gamma}$ includes more homogeneous samples to reduce variance, but risks introducing bias from heterogeneous users. We further discuss strategies for selecting $\hat{\gamma}$ in Section 4.3. Once all users in $\mathcal{V}$ have been checked, the graph is updated to $\mathcal{G}_{\hat{\gamma}}$, reflecting these new connections.
Next, for each user $u$, the algorithm aggregates data from the user and its neighbors in $\mathcal{G}_{\hat{\gamma}}$ to form a new dataset $\tilde{\mathcal{D}}_{u}$ (Line 9), which represents the samples from users who are likely to have the same preference vector as the test user. The algorithm then computes the necessary statistics for the decision phase. These statistics, similar to those defined in Equation 1, encompass a broader set of users with similar preferences as $u_{\text{test}}$, enhancing the algorithm’s decision-making capability.
Decision Phase. Finally, the algorithm selects the action $\bm{a}_{\textnormal{test}}$ for the test user $u_{\textnormal{test}}$ based on the computed statistics $\tilde{M}_{u_{\textnormal{test}}}$ and $\tilde{\bm{\theta}}_{u_{\textnormal{test}}}$ (Line 13), following a pessimistic estimate of the preference vector $\bm{\theta}_{u_{\textnormal{test}}}$ (Jin et al., 2021; Rashidinejad et al., 2021; Li et al., 2022). This approach ensures robust decision-making with offline data.
4.2 Upper Bound for Off-C2LUB
We analyze the upper bound for Algorithm 1’s suboptimality. Detailed proofs can be found in Appendices A.1. We denote the total number of samples for a set of users $\mathcal{S}$ as $N^{\mathcal{S}}=\sum_{v\in\mathcal{S}}N_{v}$. Since our analysis focuses on the suboptimality for the test user $u_{\text{test}}$, we refer to homogeneous neighbors as those neighbors of $u_{\text{test}}$ in the final graph $\mathcal{G}_{\hat{\gamma}}$ that belong to the same cluster, and to heterogeneous neighbors as those in different clusters. Moreover, in order to formalize the separation between clusters, we define the minimum heterogeneity gap between users from different clusters (Gentile et al., 2014, 2017; Wang et al., 2025):
Definition 4.1 (Minimum Heterogeneity Gap).
The preference vectors of users from different clusters are separated by at least a gap of $\gamma$. Specifically, two users $u$ and $v$ belong to the same cluster if and only if $\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}<\gamma$.
Theorem 4.2.
Let $K_{u}=\sqrt{\frac{4\log\left(\frac{2U}{\delta}\right)}{\lambda_{a}N_{u}}}$, which represents the extent of data insufficiency for user $u$. Then for given inputs $\alpha\geq 1$, $\lambda$ and $\delta$ such that $\lambda\leq d\log\left(1+\min_{u}\{N_{u}\}/(\lambda d)\right)+2\log\left(2U/\delta\right)$ and $\delta\leq\frac{2U}{1+\max_{u}\{N_{u}\}/(\lambda d)}$, the suboptimality gap for Off-C2LUB satisfies:
[…]
with probability at least $1-2\delta$, where $\tilde{O}$ omits logarithmic terms and
$\mathcal{R}_{\hat{\gamma}}(u)$ =$=\begin{cases}\{u\}\cup\left\{v\,|\,\bm{\theta}_{u}=\bm{\theta}_{v}\text{ and }K_{u}+K_{v}\leq\alpha_{r}\hat{\gamma}\text{ and }N_{v}\geq N_{\min}\right\},&\text{if }N_{u}\geq N_{\min},\\ \{u\},&\text{otherwise},\end{cases}$
$\mathcal{W}_{\hat{\gamma}}(u)$ =$=\begin{cases}\left\{v\,|\,\gamma\leq\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}\leq\hat{\gamma}\text{ and }K_{u}+K_{v}\leq\alpha_{w}\varepsilon\text{ and }N_{v}\geq N_{\min}\right\},&\text{if }N_{u}\geq N_{\min},\\ \emptyset,&\text{otherwise},\end{cases}$
$\mathcal{V}_{\hat{\gamma}}(u)$ $=\mathcal{R}_{\hat{\gamma}}(u)\cup\mathcal{W}_{\hat{\gamma}}(u),\ \eta_{\mathcal{W}_{\hat{\gamma}}(u)}=\frac{\left|\mathcal{W}_{\hat{\gamma}}(u)\right|\lambda+N^{\mathcal{W}_{\hat{\gamma}}(u)}}{|\mathcal{V}_{\hat{\gamma}}(u)|\lambda+N^{\mathcal{V}_{\hat{\gamma}}(u)}}$
for $N_{\min}=\frac{16}{\lambda_{a}^{2}}\log\left(\frac{8dU}{\lambda_{a}^{2}\delta}\right),\alpha_{r}=\frac{1}{2(\alpha+1)\sqrt{\max\{2,d\}}}$, $\alpha_{w}=\frac{1}{\alpha-1}$ and $\varepsilon=\hat{\gamma}-\gamma$.
Remark 4.3 (Notation Interpretations).
In theoretical analysis, the exact neighbors of $u_{\text{test}}$ are difficult to determine due to limited access to the full dataset $\mathcal{D}$. To address this, Algorithm 1 estimates neighbor sets using conditions from Line 5. The set $\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})$ serves as a subset estimation of homogeneous neighbors, conservatively underestimating their number to ensure that the first term in Equation 3 safely bounds bias from limited homogeneous samples. Conversely, $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$ is a superset estimation of heterogeneous neighbors, deliberately overestimated to ensure the second term captures worst-case heterogeneity via a larger $\eta_{\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})}$. These constructions allow us to derive a high-probability upper bound that remains valid despite cluster uncertainty. According to Equation 2, connections are formed only when both users meet a minimal data requirement $N_{\min}$, ensuring sufficient coverage of $\mathbb{R}^{d}$ per Assumption 3.1. Otherwise, the algorithm falls back on using only test user data. The conditions on $K_{u}+K_{v}$ for both $\mathcal{R}_{\hat{\gamma}}$ and $\mathcal{W}_{\hat{\gamma}}$ reflect our principle that only users with enough data and reliable estimates should influence decisions. A larger $\hat{\gamma}$ relaxes these constraints, potentially adding useful homogeneous samples but increasing the risk of heterogeneous bias. Finally, the constraint $\gamma\leq\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}\leq\hat{\gamma}$ ensures $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$ includes only similar users, with larger $\hat{\gamma}$ risking more bias.
Remark 4.4 (Analysis of Theorem 4.2).
The terms in Theorem 4.2 reflect both homogeneous error and heterogeneous error. The first term represents the homogeneous error associated with homogeneous neighbors in $\mathcal{R}_{\hat{\gamma}}(u_{\textnormal{test}})$. Specifically, this term decreases as the total number of samples in $\mathcal{R}_{\hat{\gamma}}(u_{\textnormal{test}})$ increases, which can be achieved by selecting a larger $\hat{\gamma}$. The second term represents the heterogeneous error, which is directly proportional to $\hat{\gamma}$. As discussed in the previous remark, increasing $\hat{\gamma}$ enlarges the heterogeneous neighbor set $\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})$ and increases $\eta_{\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})}$. Therefore, controlling $\hat{\gamma}$ introduces a trade-off between expanding $\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})$ to include more useful samples and limiting $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$ to reduce bias in the estimation, and is critical to prevent large heterogeneous errors while still leveraging enough data for prediction to decrease homogeneous errors.
4.3 Strategies for Choosing $\hat{\gamma}$
As demonstrated in Theorem 4.2, the performance of Algorithm 1 critically depends on the choice of $\hat{\gamma}$. A carefully chosen $\hat{\gamma}$ balances homogeneous and heterogeneous errors, directly affecting the algorithm’s suboptimality. In this subsection, we outline strategies for selecting $\hat{\gamma}$ under different scenarios. Figure 1 shows the influence of different choices for $\hat{\gamma}$.
Figure 1: Influence of Different $\hat{\gamma}$.
Case 1: Exact Information of $\gamma$. We set $\hat{\gamma}=\gamma$ when the exact value of $\gamma$ is known. This ensures $\varepsilon=0$ in Theorem 4.2, yielding only users from the same true cluster as $u_{\textnormal{test}}$ are included, i.e. $\eta_{\mathcal{W}_{\gamma}(u_{\textnormal{test}})}=0$. The resulting suboptimality is formalized in Corollary 4.5.
Corollary 4.5 (Theorem 4.2 for Accurate $\gamma$).
With the same notations as Theorem 4.2, if $\hat{\gamma}=\gamma$ in Algorithm 1, the suboptimality gap satisfies:
[…]
Remark 4.6 (Analysis of Corollary 4.5).
Corollary 4.5 implies that Algorithm 1 achieves suboptimality only with homogeneous error. This is because no users from outside $u_{\textnormal{test}}$’s cluster are included, eliminating heterogeneous error entirely. Furthermore, Algorithm 1 only connects homogeneous users based on similar preference vectors, avoiding the biased inclusion of heterogeneous users.
Case 2: No Prior Information on $\gamma$.
The more general and challenging case arises when no prior knowledge of $\gamma$ is available, requiring $\hat{\gamma}$ to be estimated directly from data. To guide this estimation, we define the following statistics:
$\Gamma(u,v)$ $=\left\|\hat{\bm{\theta}}_{u}-\hat{\bm{\theta}}_{v}\right\|_{2}-\alpha(\text{CI}_{u}+\text{CI}_{v}),\ M(u)=\{v\in\mathcal{V}\setminus\{u\}\mid\Gamma(u,v)>0\}$ (4)
where $\text{CI}_{u}$ denotes the confidence bound for user $u$’s estimated preference vector. Since $\Gamma(u,v)\leq\left\|\bm{\theta}_{u}-\bm{\theta}_{v}\right\|_{2}$ for $\alpha\geq 1$, it serves as a lower bound on the true difference between users’ preference vectors. If $\Gamma(u_{1},u_{2})\leq 0$, it suggests either the users belong to the same cluster or the available data is insufficient to distinguish them. Therefore, the set $M(u)$ includes users that are likely to be from different clusters than $u$, while those outside $M(u)$ either share the same preferences or cannot be confidently separated due to limited data.
Below, we propose two policies for selecting $\hat{\gamma}$ based on the users in $M(u_{\textnormal{test}})$. Note that in cases where $M(u_{\textnormal{test}})$ is empty, it is natural to choose $\hat{\gamma}=0$, as no sufficient data exists to identify any property of clusters. By selecting $\hat{\gamma}=0$, Algorithm 1 degenerates to LinUCB (Abbasi-Yadkori et al., 2011), using only the samples from $\mathcal{D}_{u_{\textnormal{test}}}$.
Underestimation Policy. The underestimation policy for choosing $\hat{\gamma}$ is defined as:
$\textstyle\hat{\gamma}=\mathbb{I}\{M(u_{\textnormal{test}})\neq\emptyset\}\cdot\min_{v\in M(u_{\textnormal{test}})}\Gamma(u_{\textnormal{test}},v).$ (5)
The intuition behind this approach stems from the property that $\Gamma(u,v)\leq\left\|\bm{\theta}_{u}-\bm{\theta}_{v}\right\|_{2}$. Since all users in $M(u_{\textnormal{test}})$ are from clusters different from that of $u_{\textnormal{test}}$, the underestimation ensures that $\hat{\gamma}$ is sufficiently small, thus avoiding the inclusion of too many biased samples in decision-making. This reduces the risk of introducing a high heterogeneous error, which explains why the minimum value of $\Gamma(u_{\textnormal{test}},v)$ is chosen from all users in $M(u_{\textnormal{test}})$. To formalize this, we present Lemma 4.7, which demonstrates how the underestimation policy can effectively restrict the size of $\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})$.
Lemma 4.7 (Effect of Underestimation).
For $\hat{\gamma}$ chosen according to Equation 5 and $\alpha\geq 1$, any user $v$ within the set for estimating heterogeneous neighbors $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$ in Theorem 4.2 also satisfies that $K_{u_{\text{test}}}+K_{v}\geq\alpha_{r}\left\|\bm{\theta}_{u_{\text{test}}}-\bm{\theta}_{v}\right\|_{2}$ for $\alpha_{r}$ defined in Theorem 4.2.
Lemma 4.7 indicates that choosing $\hat{\gamma}$ as specified in Equation 5 effectively limits the size of $\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})$. The condition $K_{u_{\textnormal{test}}}+K_{v}\geq\alpha_{r}\|\bm{\theta}_{u_{\textnormal{test}}}-\bm{\theta}_{v}\|_{2}$ ensures that only heterogeneous users without much data are included in $\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})$, further reducing the impact of biased samples. However, while this policy is effective at restricting $\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})$, the small value of $\hat{\gamma}$ also limits the size of $\mathcal{R}_{\hat{\gamma}}(u_{\textnormal{test}})$, the set of users within the same cluster as $u_{\textnormal{test}}$. This reduction in sample size may affect the quality of estimation and increase the homogeneous error. Therefore, the underestimation policy is most suitable in scenarios where avoiding biased samples and minimizing heterogeneous error are critical, even at the expense of reducing the number of available samples within the same cluster. For example, in medical treatment scenarios, introducing heterogeneous error could pose significant risks to patients, making it critical to prioritize avoiding such errors over utilizing more available samples.
Overestimation Policy. In contrast to underestimation, the overestimation policy uses an upper-bound estimate to choose $\hat{\gamma}$. The policy is defined as:
$\textstyle\hat{\gamma}=\mathbb{I}\{M(u_{\textnormal{test}})\neq\emptyset\}\cdot\min_{v\in M(u_{\textnormal{test}})}\tilde{\Gamma}(u_{\textnormal{test}},v),$ (6)
where $\tilde{\Gamma}(u_{\textnormal{test}},v)=\left\|\hat{\bm{\theta}}_{u_{\textnormal{test}}}-\hat{\bm{\theta}}_{v}\right\|_{2}+\alpha(\text{CI}_{u_{\textnormal{test}}}+\text{CI}_{v})$ is an upper-bound estimate of the gap between the test user $u_{\textnormal{test}}$ and another user $v$.
Lemma 4.8 (Effect of Overestimation).
For $\hat{\gamma}$ chosen as in Equation 6, if $M(u_{\text{test}})\neq\emptyset$, $\hat{\gamma}\geq\gamma$.
By choosing $\hat{\gamma}\geq\gamma$ according to Equation 6, Algorithm 1 ensures that samples from homogeneous users with sufficient data are included in $\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})$. This significantly reduces homogeneous error by leveraging a larger dataset. However, as $\hat{\gamma}\geq\gamma$, this policy may also introduce some biased users into $\mathcal{W}_{\hat{\gamma}}(u_{\textnormal{test}})$, increasing the heterogeneous error. The trade-off here is between reducing homogeneous error and potentially increasing heterogeneous error. Therefore, the overestimation policy is particularly effective in scenarios where combining a large amount of data from diverse users is critical, e.g., for high-dimensional preference vectors, and the goal is to minimize homogeneous error rather than strictly controlling heterogeneous error.
5 Algorithm for Sufficient Dataset: Offline - Clustering of Bandits
When $\gamma$ is unknown, Algorithm 1 (Off-C2LUB) faces a trade-off: setting $\hat{\gamma}$ too small increases homogeneous error, while setting it too large introduces heterogeneous error, as discussed in Section 4.3. To address this, we propose a second algorithm, Offline Clustering of Bandits (Off-CLUB), that avoids explicitly estimating $\gamma$ and achieves strong performance when the dataset is sufficiently large.
5.1 Algorithm 2: Off-CLUB
Off-CLUB shares the same two-phase structure as Off-C2LUB, consisting of a Cluster Phase and a Decision Phase. Due to space constraints, full details and pseudo-code (Algorithm 2) are deferred to Appendix D. The main difference lies in graph construction: Off-CLUB starts with a complete graph and removes edges between users with dissimilar preferences, whereas Off-C2LUB begins with an empty graph and adds edges only when sufficient similarity is detected.
The algorithm takes similar inputs to Algorithm 1 but does not require $\hat{\gamma}$. During the Cluster Phase, it iterates through all edges $(u_{1},u_{2})\in\mathcal{E}$ and removes those that satisfying:
$\textstyle\left\|\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}\right\|_{2}>\alpha\left(\text{CI}_{u_{1}}+\text{CI}_{u_{2}}\right),$ (7)
indicating that users $u_{1}$ and $u_{2}$ likely belong to different clusters. After edge removal, the resulting graph $\tilde{\mathcal{G}}$ correctly clusters users with sharing preference vectors, provided that all users have sufficient samples as discussed in Section 5.2. Finally, the subsequent steps, including data aggregation, computation of statistics, and the Decision Phase, are largely similar to those in Off-C2LUB.
5.2 Upper Bound for Off-CLUB under Sufficient Data
The effectiveness of Off-CLUB depends on correctly removing edges between users in different clusters, which requires sufficiently small confidence bounds $\text{CI}_{u}$. This condition is typically met when each user has a large number of samples. Conversely, users with limited samples may remain incorrectly connected, causing users from different clusters to be grouped together. Thus, to formalize the conditions for Off-CLUB’s success, we introduce the sufficient data assumption:
Assumption 5.1 ($\delta$-Sufficient Dataset).
The dataset $\mathcal{D}_{u}$ is called $\delta$-sufficient for user $u$ if $N_{u}\geq\max\left\{\frac{16}{\lambda_{a}^{2}}\log\left(\frac{8dU}{\lambda_{a}^{2}\delta}\right),\,\frac{512d}{\gamma^{2}\lambda_{a}}\log\left(\frac{2U}{\delta}\right)\right\}$ for some $\delta\in(0,1)$. Furthermore, we assume that the full dataset $\mathcal{D}=\{\mathcal{D}_{u}\}_{u=1}^{U}$ is completely $\delta$-sufficient, i.e., each user’s dataset $\mathcal{D}_{u}$ is $\delta$-sufficient.
This assumption ensures that each user has enough samples to accurately estimate preference vectors, enabling Off-CLUB to effectively cluster users. Theorem 5.2 (proof in Appendix A.4) establishes the performance guarantees of Off-CLUB under this assumption.
Theorem 5.2 (Upper Bound for Off-CLUB).
For any $\delta\in(0,\frac{1}{2})$, assume $\mathcal{D}$ is completely $\delta$-sufficient. For a given input $\alpha=1$ and $\lambda>0$, with probability at least $1-2\delta$, the suboptimality gap satisfies
$\textnormal{SubOpt}(u_{\textnormal{test}},\textnormal{Off-CLUB},A_{\textnormal{test}})\leq\tilde{O}\left(\sqrt{\frac{d}{N^{\mathcal{V}(j_{u_{\textnormal{test}}})}}}\right)$
where $N^{\mathcal{V}(j_{u_{\textnormal{test}}})}=\sum_{v\in\mathcal{V}(j_{u_{\textnormal{test}}})}N_{v}$ denotes the total samples from users in $u_{\text{test}}$’s cluster.
Remark 5.3 (Suboptimality Analysis).
Here $N^{\mathcal{V}(j_{u_{\textnormal{test}}})}$ depends on both the number of users in the cluster and the samples per user. Theorem 5.2 matches the lower bound discussed in Section 6. A key requirement for achieving this bound is the sufficiency of data for each user (5.1). When this assumption is violated, Off-CLUB may incorporate samples from highly biased users with preference vectors largely different with $u_{\text{test}}$, degrading performance. In contrast, as shown in Theorem 4.2, Off-C2LUB mitigates this issue by restricting heterogeneous neighbors to a much smaller set $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$, controlled by the choice of $\hat{\gamma}$. Thus, while Off-CLUB performs well with sufficient data, Off-C2LUB remains a more robust algorithm in general scenarios where Assumption 5.1 does not hold. We validate this intuition with experiments in Section 7.
6 Lower Bound for Off-ClusBand and Optimality
This section presents the theoretical lower bound on suboptimality for the Off-ClusBand problem. Theorem 6.1 establishes the bound under minimal data and regularity conditions, with the proof in Appendix A.5.
Theorem 6.1 (Off-ClusBand Lower Bound).
[…]
$\inf_{\pi}\sup_{\mathcal{Q}\in\mathcal{I}_{u_{\textnormal{test}}}}\mathbb{E}[\textnormal{SubOpt}_{\mathcal{Q}}(u_{\textnormal{test}},\pi,A_{\textnormal{test}})]\geq\sqrt{\frac{8d}{N^{\mathcal{V}(j_{u_{\textnormal{test}}})}}}.$
Remark 6.2 (Comparison under Sufficient Data).
The suboptimality bound in Theorem 5.2, $\tilde{O}\left(\sqrt{\frac{d}{N^{\mathcal{V}(j_{u_{\text{test}}})}}}\right)$, matches the lower bound in Theorem 6.1 up to logarithmic factors, proving that Algorithm 2 (Off-CLUB) achieves near-optimal performance when the dataset satisfies Assumption 5.1. Similarly, Algorithm 1 (Off-C2LUB) with $\hat{\gamma}=\gamma$ also meets the lower bound under the same assumption with the input $\alpha=1$. Consequently, all users in the same cluster as $u_{\text{test}}$ are correctly identified, guaranteeing optimal sample utilization.
Remark 6.3 (Comparison under Limited Data).
When data is insufficient, Algorithm 2 (Off-CLUB) often clusters heterogeneous users together, incorporating biased samples and degrading decision quality. In contrast, Algorithm 1 (Off-C2LUB) with $\hat{\gamma}=\gamma$ mitigates this issue by excluding biased samples, as shown in Corollary 4.5. Yet, limited data prevents it from identifying all homogeneous users correctly of the test user, leading to a minor gap between its upper bound and the theoretical lower bound in Theorem 6.1. Moreover, as shown in Theorem 4.2, when $\gamma$ is unknown, Algorithm 1 restricts the heterogeneous neighbors of $u_{\text{test}}$ to the set $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$, which includes only those users whose distance to $u_{\text{test}}$ is at most $\hat{\gamma}$. In contrast, Algorithm 2 may retain connections between highly heterogeneous users due to insufficient data, increasing bias.
7 Experiments
Experimental Setup. We consider $U=1k$ users, partitioned into $J=10$ clusters, with datasets ranging from $5k$ to $1M$ points, equally split between offline learning and evaluation. The reward model is $r_{u}^{i}=\langle\bm{\theta}_{u},\bm{a}_{u}^{i}\rangle+\epsilon_{u}^{i}$, where $\bm{a}_{u}^{i}$ and $\bm{\theta}_{u}$ are item and user features, and $\epsilon_{u}^{i}$ is 1-sub-Gaussian noise. We benchmark against LinUCB_IND, CLUB, an enhanced DBSCAN variant (DBSCAN_Improve), and an improved variant of XMeans (Pelleg et al., 2000) (XMeans_Improve). We use both a synthetic dataset, consisting of $d=20$-dimensional Gaussian vectors, and two real-world datasets (Yelp [^1] https://business.yelp.com/data/resources/open-dataset/. and MovieLens Harper and Konstan (2015)), sampled from the top 1,000 users and items. User preference vectors are extracted using Singular Value Decomposition (SVD). Offline data is generated using random or LinUCB-based item selection. More details can be found in Appendix C.
Performance with Insufficient Dataset. Figure 2 shows the average suboptimality gap under different user distributions (equal and semi-random) and dataset sizes ($|\mathcal{D}|$). Each subplot represents a unique user distribution and dataset. For Off-C2LUB, $\hat{\gamma}$ is optimized in an equal distribution scenario and applied to the semi-random distribution. When $|\mathcal{D}|\in[20k,100k]$, Off-C2LUB_Optimistic (Pessimistic) significantly reduces the suboptimality gap. On the synthetic dataset, it outperforms Off-CLUB by 33.0% (31.7%) and other baselines by at least 84.7% (85.9%). On Yelp, it reduces the gap by 55.9% (58.8%) over Off-CLUB and at least 73.1% (74.8%) over baselines. On MovieLens, it reduces the gap by 41.3% (44.9%) over Off-CLUB and at least 66.0% (68.1%) over baselines.
Performance with Sufficient Dataset. Table 3 shows that Off-CLUB performs better as $|\mathcal{D}|$ increases and it is more likely to have sufficient data. Furthermore, as $|\mathcal{D}|$ increases, both Off-CLUB and Off-C2LUB consistently maintain superior performance, achieving the lowest suboptimality gaps.
Improvement with Estimated $\hat{\gamma}$. We evaluate various strategies under insufficient data using synthetic and Yelp datasets (Figure 3), with $|\mathcal{D}|=30k$; additional results are in Appendix C. Solid lines indicate methods that use $\hat{\gamma}$, dashed lines those that do not. On the synthetic dataset, relative suboptimality gaps compared to the optimal $\hat{\gamma}$ of Off-C2LUB are: DBSCAN (5.90×5.90$\times$5.90 ×), CLUB (2.55×2.$5$), Off-C2LUB_Pessimistic (1.50×1.50$\times$1.50 ×), and Off-C2LUB_Optimistic (1.22×1.$2$). Similar trends hold for Yelp: DBSCAN (3.10×3.10$\times$3.10 ×), CLUB (2.26×2.26$\times$2.26 ×), Pessimistic (1.41×1.41$\times$1.41 ×), and Optimistic (1.24×1.24$\times$1.24 ×).
Figure 2: Suboptimality Gap under Varying User Distributions with Limited Data
Figure 3: Suboptimality Gap under Varying $\hat{\gamma}$
Table 2: Effect of Data Size on Suboptimality
Algorithm Dataset size $0.2M$ $0.4M$ $0.6M$ $0.8M$ $1M$
DBSCAN_Improve 0.130149 0.058302 0.024758 0.011941 0.006520
XMeans_Improve 0.129364 0.058108 0.024791 0.011955 0.006536
CLUB 0.013153 0.000019 0.000012 0.000009 0.000007
Off-CLUB 0.000038 0.000018 0.000012 0.000009 0.000007
LinUCB 0.140748 0.140993 0.140925 0.140898 0.140844
LinUCB_IND 0.144851 0.066494 0.028189 0.013586 0.007561
SCLUB 0.180909 0.131105 0.050664 0.051058 0.051328
Off-C2LUB_Pessimistic 0.000037 0.000019 0.000013 0.000011 0.000008
Off-C2LUB_Optimistic 0.000041 0.000018 0.000012 0.000009 0.000007
8 Conclusion
In this paper, we address the offline clustering of bandits (Off-ClusBand) problem, proposing two algorithms: Off-C2LUB, effective in general scenarios, and Off-CLUB, which excels with sufficient data. We analyze the influence of $\hat{\gamma}$ on Off-C2LUB, offering policies for cases where $\gamma$ is known or unknown. Additionally, we establish a theoretical lower bound for Off-ClusBand. Experiments on synthetic and real-world datasets validate the robustness and efficiency of our methods. Future work may explore more general reward structures beyond linear models that this paper does not cover, such as nonlinear or nonparametric formulations. It is also promising to consider hybrid settings that combine offline data with online interaction for improved adaptability in real-world scenarios.
Appendix A Detailed Proofs
A.1 Proof of Theorem 4.2
To prove Theorem 4.2, we first introduce several key notations and supporting lemmas, which we present and prove sequentially.
Table 4: Notations Defined in Proofs
Notation Description
$V_{\hat{\gamma}}(u)$ The set consisting of user $u$ and its neighbors in the graph $\mathcal{G}_{\hat{\gamma}}$.
$R_{\hat{\gamma}}(u)$ Users in $V_{\hat{\gamma}}(u)$ that share the same preference vector as user $u$.
$W_{\hat{\gamma}}(u)$ Users in $V_{\hat{\gamma}}(u)$ that do not share the same preference vector as user $u$.
$W_{\hat{\gamma}}^{j}(u)$ Users in $W_{\hat{\gamma}}(u)$ who belong to cluster $V(j)$.
$J_{\hat{\gamma}}(u)$ Indices of distinct clusters represented in $W_{\hat{\gamma}}(u)$.
To characterize the contributions of different types of neighbors in the estimation process, we define the weighted fraction of samples from $R_{\hat{\gamma}}(u)$ relative to $V_{\hat{\gamma}}(u)$ as:
$\eta_{R_{\hat{\gamma}}(u)}=\frac{|R_{\hat{\gamma}}(u)|\lambda+N^{R_{\hat{\gamma}}(u)}}{|V_{\hat{\gamma}}(u)|\lambda+N^{V_{\hat{\gamma}}(u)}}.$
Here, $\lambda$ is the regularization parameter used in constructing each user’s matrix $M_{u}$. Analogous definitions apply to $\eta_{W_{\hat{\gamma}}(u)}$ and $\eta_{W_{\hat{\gamma}}^{j}(u)}$, and by construction, we have:
$\eta_{R_{\hat{\gamma}}(u)}+\eta_{W_{\hat{\gamma}}(u)}=1,$
since $V_{\hat{\gamma}}(u)=R_{\hat{\gamma}}(u)\cup W_{\hat{\gamma}}(u)$ and the two sets are disjoint.
Lemma A.1.
With the notations defined in Table 4, the suboptimality gap of Off-C2LUB satisfies:
$\textnormal{SubOpt}(u_{\textnormal{test}},\textnormal{Off-C}^{2}\textnormal{LUB},A_{\textnormal{test}})\leq\frac{(2\sqrt{2}+1)\beta(V_{\hat{\gamma}}(u_{\text{test}}),\delta)}{\sqrt{\lambda_{a}N^{V_{\hat{\gamma}}(u_{\text{test}})}}}+\frac{4\lambda U}{\lambda_{a}N^{V_{\hat{\gamma}}(u_{\text{test}})}}+\frac{4\eta_{W_{\hat{\gamma}}(u_{\text{test}})}}{\lambda_{a}}\cdot\hat{\gamma}$
with probability at least $1-2\delta$, where
$\beta(\mathcal{S},\delta)=\sqrt{d\log\left(1+\frac{N^{\mathcal{S}}}{|\mathcal{S}|\lambda d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{|\mathcal{S}|\lambda}$
for any set $\mathcal{S}$.
Lemma A.2.
The bound in Lemma A.1 can be simplified as:
[…]
Lemma A.3.
For the notations and inputs in Theorem 4.2, the following containment relations hold:
$\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})\subseteq R_{\hat{\gamma}}(u_{\text{test}}),\quad\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})\supseteq W_{\hat{\gamma}}(u_{\text{test}}),\quad\eta_{\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})}\geq\eta_{W_{\hat{\gamma}}(u_{\text{test}})}.$
Proof of Lemma A.1.
Let $\beta(\mathcal{S},\delta)=\sqrt{d\log\left(1+\frac{\sum_{u\in\mathcal{S}}N_{u}}{|\mathcal{S}|\lambda d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{|\mathcal{S}|\lambda}$ for any set $\mathcal{S}$. We begin by defining the following confidence event for each user $u$:
$\mathcal{F}_{u}^{\prime}(\delta)=\left\{\left\|\bm{\theta}_{u}-\hat{\bm{\theta}}_{u}\right\|_{M_{u}}\leq\beta(\{u\},\delta)\right\},$
where
$M^{\mathcal{S}}=\lambda|\mathcal{S}|\cdot I+\sum_{u\in\mathcal{S}}\sum_{i=1}^{N_{u}}\bm{a}_{u}^{i}(\bm{a}_{u}^{i})^{\top},\quad\bm{b}^{\mathcal{S}}=\sum_{u\in\mathcal{S}}\sum_{i=1}^{N_{u}}r_{u}^{i}\bm{a}_{u}^{i},\quad\bm{\theta}^{\mathcal{S}}=(M^{\mathcal{S}})^{-1}\bm{b}^{\mathcal{S}}.$
By Lemma B.1, we have $\mathbb{P}[\mathcal{F}_{v}^{\prime}(\delta)]\geq 1-\frac{\delta}{2U}$ for each $v\in\mathcal{U}$. Therefore, applying the union bound over all users yields:
$\mathbb{P}\left[\bigcap_{v\in\mathcal{U}}\mathcal{F}_{v}^{\prime}(\delta)\right]\geq 1-\frac{\delta}{2}.$
Furthermore, for both $R_{\hat{\gamma}}(u)$ and $W_{\hat{\gamma}}(u)$, the clustering condition in Equation 2 requires that a user $v$ can be included in either set only if both $N_{u}$ and $N_{v}$ are at least $\frac{16}{\lambda_{a}^{2}}\log\left(\frac{8dU}{\lambda_{a}^{2}\delta}\right)$. Hence, the case where $N_{u}<\frac{16}{\lambda_{a}^{2}}\log\left(\frac{8dU}{\lambda_{a}^{2}\delta}\right)$ is trivial since the problem can be simplified to an offline linear contextual bandit problem in [Li et al., 2022].
For the non-trivial case where $N_{u}\geq\frac{16}{\lambda_{a}^{2}}\log\left(\frac{8dU}{\lambda_{a}^{2}\delta}\right)$, we apply Lemma B.2 and the union bound to obtain that for any $v\in R_{\hat{\gamma}}(u)\cup W_{\hat{\gamma}}(u)$, with probability at least $1-\frac{3}{2}\delta$, the following inequality holds:
$\left\|\bm{\theta}_{v}-\hat{\bm{\theta}}_{v}\right\|_{2}\leq\frac{\left\|\bm{\theta}_{v}-\hat{\bm{\theta}}_{v}\right\|_{M_{v}}}{\sqrt{\lambda_{\min}(M_{v})}}\leq\frac{\beta(\{v\},\delta)}{\sqrt{\lambda+\lambda_{a}N_{v}/2}}\leq\text{CI}_{v}.$
Here, since $\hat{\gamma}\geq 0$, for any pair of connected users $(u_{1},u_{2})$, we have:
$\|\bm{\theta}_{u_{1}}-\bm{\theta}_{u_{2}}\|_{2}$ $=\|\bm{\theta}_{u_{1}}-\hat{\bm{\theta}}_{u_{1}}+\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}+\hat{\bm{\theta}}_{u_{2}}-\bm{\theta}_{u_{2}}\|_{2}$
$\leq\|\bm{\theta}_{u_{1}}-\hat{\bm{\theta}}_{u_{1}}\|_{2}+\|\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}\|_{2}+\|\hat{\bm{\theta}}_{u_{2}}-\bm{\theta}_{u_{2}}\|_{2}$
$\leq\text{CI}_{u_{1}}+\|\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}\|_{2}+\text{CI}_{u_{2}}\leq\hat{\gamma},$
where the last inequality follows from the connection condition in Algorithm 1. Therefore, $R_{\hat{\gamma}}(u)$ includes users that share the same preference vector as user $u$, while $W_{\hat{\gamma}}(u)$ includes users whose preference vectors differ from that of $u$ by at most $\hat{\gamma}=\gamma+\varepsilon$.
We now turn to bounding the suboptimality term $\text{SubOpt}(u,\pi,A_{\text{test}})=\left\langle\bm{\theta}_{u},\bm{a}_{u}^{*}\right\rangle-\left\langle\bm{\theta}_{u},\bm{a}_{\text{test}}\right\rangle$. For notational simplicity, we let:
$V=V_{\hat{\gamma}}(u),\quad R=R_{\hat{\gamma}}(u),\quad W=W_{\hat{\gamma}}(u),\quad W^{j}=W_{\hat{\gamma}}^{j}(u),\quad\tilde{J}=J_{\hat{\gamma}}(u).$
Moreover, for a user set $\mathcal{S}$, we define the following aggregated statistics:
$M^{\mathcal{S}}=\lambda|\mathcal{S}|I+\sum_{u\in\mathcal{S}}\sum_{i=1}^{N_{u}}\bm{a}_{u}^{i}(\bm{a}_{u}^{i})^{\top},\quad\bm{b}^{\mathcal{S}}=\sum_{u\in\mathcal{S}}\sum_{i=1}^{N_{u}}r_{u}^{i}\bm{a}_{u}^{i},\quad\hat{\bm{\theta}}^{\mathcal{S}}=(M^{\mathcal{S}})^{-1}\bm{b}^{\mathcal{S}}.$
Recall that $\bm{b}^{V}=\bm{b}^{R}+\sum_{j\in\tilde{J}}\bm{b}^{W^{j}}$ and $M^{V}=M^{R}+\sum_{j\in\tilde{J}}M^{W^{j}}$, we can decompose $\hat{\bm{\theta}}^{V}$ as:
$\hat{\bm{\theta}}^{V}$ $=(M^{V})^{-1}\bm{b}^{V}=(M^{V})^{-1}(\bm{b}^{R}+\sum_{j\in\tilde{J}}\bm{b}^{W^{j}})$
$=(M^{V})^{-1}M^{R}(M^{R})^{-1}\bm{b}^{R}+\sum_{j\in\tilde{J}}(M^{V})^{-1}M^{W^{j}}(M^{W^{j}})^{-1}\bm{b}^{W^{j}}$
$=(M^{V})^{-1}M^{R}\hat{\bm{\theta}}^{R}+\sum_{j\in\tilde{J}}(M^{V})^{-1}M^{W^{j}}\hat{\bm{\theta}}^{W^{j}}.$
Let $\varepsilon_{v}^{i}=r_{v}^{i}-\bm{\theta}_{v}^{\top}\bm{a}_{v}^{i}$ denote the error for the sample $(\bm{a}_{v}^{i},r_{v}^{i})\in\mathcal{D}_{v}$ where $i\in[N_{v}]$. Hence, we have:
$\left\langle\hat{\bm{\theta}}^{V}-\bm{\theta}_{u},\bm{a}_{\text{test}}\right\rangle$ $=\left\langle(M^{V})^{-1}M^{R}\hat{\bm{\theta}}^{R}+\sum_{j\in\tilde{J}}(M^{V})^{-1}M^{W^{j}}\hat{\bm{\theta}}^{W^{j}}-\bm{\theta}_{u},\bm{a}_{\text{test}}\right\rangle$ (9)
$=\left\langle(M^{V})^{-1}M^{R}(\hat{\bm{\theta}}^{R}-\bm{\theta}_{u})+\sum_{j\in\tilde{J}}(M^{V})^{-1}M^{W^{j}}(\hat{\bm{\theta}}^{W^{j}}-\bm{\theta}_{u}),\bm{a}_{\text{test}}\right\rangle$
$=\left\langle(M^{V})^{-1}M^{R}(\hat{\bm{\theta}}^{R}-\bm{\theta}_{u}),\bm{a}_{\text{test}}\right\rangle+\left\langle\sum_{j\in\tilde{J}}(M^{V})^{-1}M^{W^{j}}(\hat{\bm{\theta}}^{W^{j}}-\bm{\theta}_{u}),\bm{a}_{\text{test}}\right\rangle$
$=\left\langle(M^{V})^{-1}\left(\bm{b}^{R}-M^{R}\bm{\theta}_{u}+\sum_{j\in\tilde{J}}(\bm{b}^{W^{j}}-M^{W^{j}}\bm{\theta}^{j})\right),\bm{a}_{\text{test}}\right\rangle$
$\quad+\left\langle\sum_{j\in\tilde{J}}(M^{V})^{-1}M^{W^{j}}(\bm{\theta}^{j}-\bm{\theta}_{u}),\bm{a}_{\text{test}}\right\rangle$
$=\left\langle(M^{V})^{-1}\sum_{v\in V}\sum_{i=1}^{N_{v}}\varepsilon_{v}^{i}\bm{a}_{v}^{i},\bm{a}_{\text{test}}\right\rangle-\left\langle(M^{V})^{-1}\lambda\left(|R|\bm{\theta}_{u}+\sum_{j\in\tilde{J}}|W^{j}|\bm{\theta}^{j}\right),\bm{a}_{\text{test}}\right\rangle$
$\quad+\left\langle\sum_{j\in\tilde{J}}(M^{V})^{-1}M^{W^{j}}(\bm{\theta}^{j}-\bm{\theta}_{u}),\bm{a}_{\text{test}}\right\rangle$ (10)
Applying the Cauchy–Schwarz inequality, we obtain from Equation 10:
$\left\langle\hat{\bm{\theta}}^{V}-\bm{\theta}_{u},\bm{a}_{\text{test}}\right\rangle\leq\;$ $\left\|(M^{V})^{-1/2}\bm{a}_{\text{test}}\right\|_{2}\cdot\left\|\sum_{v\in V}\sum_{i=1}^{N_{v}}\varepsilon_{v}^{i}\bm{a}_{v}^{i}\right\|_{(M^{V})^{-1}}$
$+\left\|\lambda(M^{V})^{-1}\left(|R|\bm{\theta}_{u}+\sum_{j\in\tilde{J}}|W^{j}|\bm{\theta}^{j}\right)\right\|_{2}\cdot\|\bm{a}_{\text{test}}\|_{2}$
$+\left\|\sum_{j\in\tilde{J}}(M^{V})^{-1}M^{W^{j}}(\bm{\theta}^{j}-\bm{\theta}_{u})\right\|_{2}$ (11)
We now bound each term using matrix norm properties and known results from Equation 11:
$\left\langle\hat{\bm{\theta}}^{V}-\bm{\theta}_{u},\bm{a}_{\text{test}}\right\rangle\leq\;$ $\frac{\beta(V,\delta)}{\sqrt{\lambda_{\min}(V)}}+\frac{\lambda|V|}{\lambda_{\min}(V)}+\frac{\lambda_{\max}(W)}{\lambda_{\min}(V)}\cdot\hat{\gamma}$
[…]
with probability at least $1-2\delta$.
Here, Equation 9 follows from the decomposition $M^{V}=M^{R}+\sum_{j\in\tilde{J}}M^{W^{j}}$; Equation 10 uses the fact that $M^{\mathcal{S}}=\lambda|\mathcal{S}|I+\sum_{v\in\mathcal{S}}\sum_{i=1}^{N_{v}}\bm{a}_{v}^{i}(\bm{a}_{v}^{i})^{\top}$ and the definition of reward noise $\varepsilon_{v}^{i}$. Inequality (11) is due to the Cauchy–Schwarz inequality. The first inequality in (12) uses eigenvalue bounds and Theorem 1 in [Abbasi-Yadkori et al., 2011], while the second inequality follows from bounding eigenvalues via the number of samples and applying Lemma B.2.
Therefore, we obtain:
$-\langle\bm{\theta}_{u},\bm{a}_{\text{test}}\rangle\leq\Psi-\left\langle\hat{\bm{\theta}}^{V},\bm{a}_{\text{test}}\right\rangle.$
Recall that the selected action $\bm{a}_{\text{test}}$ satisfies the confidence bound:
$\left(\hat{\bm{\theta}}^{V}\right)^{\top}\bm{a}_{\text{test}}-\beta(V,\delta)\|\bm{a}_{\text{test}}\|_{(M^{V})^{-1}}\geq\left(\hat{\bm{\theta}}^{V}\right)^{\top}\bm{a}_{u}^{*}-\beta(V,\delta)\|\bm{a}_{u}^{*}\|_{(M^{V})^{-1}}.$
Hence, we can derive the following bound for the suboptimality:
$\text{SubOpt}(u,\pi,A_{\text{test}})$ $=\langle\bm{\theta}_{u},\bm{a}_{u}^{*}\rangle-\langle\bm{\theta}_{u},\bm{a}_{\text{test}}\rangle$
$\leq\langle\bm{\theta}_{u},\bm{a}_{u}^{*}\rangle-\left\langle\hat{\bm{\theta}}^{V},\bm{a}_{\text{test}}\right\rangle+\Psi$
$\leq\left\langle\bm{\theta}_{u}-\hat{\bm{\theta}}^{V},\bm{a}_{u}^{*}\right\rangle+\beta(V,\delta)\left(\|\bm{a}_{u}^{*}\|_{(M^{V})^{-1}}-\|\bm{a}_{\text{test}}\|_{(M^{V})^{-1}}\right)+\Psi$
$\leq 2\Psi+\frac{\sqrt{2}\beta(V,\delta)}{\sqrt{\lambda_{\min}(M^{V})}}$
$\leq 2\Psi+\frac{\beta(V,\delta)}{\sqrt{\lambda_{a}N^{V}}}$
with probability at least $1-2\delta$, which completes the proof of Lemma A.1. ∎
Proof of Lemma A.2.
From Lemma A.1 we know that
$\textnormal{SubOpt}(u_{\text{test}},\text{Off-NCLUB},A_{\text{test}})$ $\leq\tilde{O}\left(\frac{\sqrt{d}}{\sqrt{N^{{V}_{\hat{\gamma}}(u_{\text{test}})}}}+\eta_{{W}_{\hat{\gamma}}(u_{\text{test}})}\hat{\gamma}\right)$
$\leq\tilde{O}\left(\sqrt{\frac{d}{N^{R_{\hat{\gamma}}(u_{\text{test}})}}}+\eta_{{W}_{\hat{\gamma}}(u_{\text{test}})}\cdot\hat{\gamma}\right)$ (13)
which ends the proof of Lemma A.2.
Proof of Lemma A.3.
We now proceed to prove Lemma A.3.
Proof of $\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})\subseteq R_{\hat{\gamma}}(u_{\text{test}})$.
We begin by analyzing the estimation set $\mathcal{R}_{\hat{\gamma}}(u)$. To establish this subset relation, it suffices to show that, for
$\alpha_{r}=\frac{1}{2(\alpha+1)\sqrt{2\max\{2,d\}}},$
any user $v$ such that $\bm{\theta}_{u}=\bm{\theta}_{v}$, $K_{u}+K_{v}\leq\alpha_{r}\hat{\gamma}$ and $N_{v}\geq N_{\min}$ is included in $R_{\hat{\gamma}}(u)$ with high probability when $N_{u}\geq N_{\min}$.
Note that the condition $K_{u}+K_{v}<\frac{\hat{\gamma}}{2(\alpha+1)\sqrt{2\max\{2,d\}}}$ implies
$(\alpha+1)(\text{CI}_{u}+\text{CI}_{v})$ $\leq\frac{2(\alpha+1)\sqrt{4\max\{2,d\}\log(\frac{2U}{\delta})}}{\sqrt{\lambda_{a}N_{u}}}+\frac{2(\alpha+1)\sqrt{4\max\{2,d\}\log(\frac{2U}{\delta})}}{\sqrt{\lambda_{a}N_{v}}}$
$<\hat{\gamma},$
under the parameter selections for $\lambda$ and $\delta$ such that
$\lambda\leq d\log\left(1+\frac{\min_{w}\{N_{w}\}}{\lambda d}\right)+2\log\left(\frac{2U}{\delta}\right),\quad\log\left(1+\frac{\max_{w}\{N_{w}\}}{\lambda d}\right)\leq\log\left(\frac{2U}{\delta}\right).$
Given that $\bm{\theta}_{u}=\bm{\theta}_{v}$, we further observe:
$\hat{\gamma}$ $>(\alpha+1)(\text{CI}_{u}+\text{CI}_{v})$
$=\alpha(\text{CI}_{u}+\text{CI}_{v})+\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}+\text{CI}_{u}+\text{CI}_{v}$
$\geq\|\hat{\bm{\theta}}_{u}-\hat{\bm{\theta}}_{v}\|_{2}+\alpha(\text{CI}_{u}+\text{CI}_{v}),$
with probability at least $1-2\delta$, by Lemma A.1. This directly implies that the connection condition in Algorithm 1 is satisfied, and therefore user $v$ is included in $R_{\hat{\gamma}}(u)$.
Proof of $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})\supseteq W_{\hat{\gamma}}(u_{\text{test}})$.
It suffices to show that if $K_{u}+K_{v}>\frac{\varepsilon}{\alpha-1}$ and $\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}\geq\gamma$, then user $v$ is not included in $\mathcal{W}_{\hat{\gamma}}(u)$ with probability at least $1-2\delta$ when $\min\{N_{u},N_{v}\}\geq N_{\min}$.
Observe that the condition $K_{u}+K_{v}>\frac{\varepsilon}{\alpha-1}$ implies:
$(\alpha-1)(\text{CI}_{u}+\text{CI}_{v})$ $\geq(\alpha-1)\left(\sqrt{\frac{4\log\left(\frac{2U}{\delta}\right)}{\lambda_{a}N_{u}}}+\sqrt{\frac{4\log\left(\frac{2U}{\delta}\right)}{\lambda_{a}N_{v}}}\right)$
$>\varepsilon\geq\hat{\gamma}-\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}.$
Therefore, we can deduce that:
$\|\hat{\bm{\theta}}_{u}-\hat{\bm{\theta}}_{v}\|_{2}+\alpha(\text{CI}_{u}+\text{CI}_{v})$ $\geq\alpha(\text{CI}_{u}+\text{CI}_{v})+\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}-\text{CI}_{u}-\text{CI}_{v}$
$>\hat{\gamma},$
which means that the connection condition in Algorithm 1 does not hold. Hence, user $v$ will not be connected to $u$, and thus not included in $\mathcal{W}_{\hat{\gamma}}(u)$. Therefore, the superset condition holds with probability at least $1-2\delta$, provided $\alpha_{w}\in\left(0,\frac{1}{\alpha-1}\right)$.
Proof of $\eta_{\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})}\geq\eta_{W_{\hat{\gamma}}(u_{\text{test}})}$.
Since we have shown that $\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})\subseteq R_{\hat{\gamma}}(u_{\text{test}})$ and $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})\supseteq W_{\hat{\gamma}}(u_{\text{test}})$, the denominator in the definition of $\eta_{\mathcal{W}_{\hat{\gamma}}(u)}$ is larger while the numerator is at least as large, leading to:
$\eta_{\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})}\geq\eta_{W_{\hat{\gamma}}(u_{\text{test}})}.$
This concludes the proof of Lemma A.3.
We are now ready to prove Theorem 4.2. From Lemma A.2, we know that the suboptimality gap increases with the size of $W_{\hat{\gamma}}(u_{\text{test}})$ and decreases with the size of $R_{\hat{\gamma}}(u_{\text{test}})$. Therefore, by conservatively replacing $R_{\hat{\gamma}}(u_{\text{test}})$ with its subset estimate $\mathcal{R}_{\hat{\gamma}}(u_{\text{test}})$, and replacing $W_{\hat{\gamma}}(u_{\text{test}})$ with its superset estimate $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$, we can ensure a valid high-probability upper bound on the suboptimality gap.
A.2 Proof of Lemma 4.7
Proof of Lemma 4.7.
Recall that we define
$\hat{\gamma}=\mathbb{I}\{M(u_{\textnormal{test}})\neq\emptyset\}\cdot\min_{v\in M(u_{\textnormal{test}})}\Gamma(u_{\textnormal{test}},v),$ (14)
where $M(u_{\textnormal{test}})$ is the set of users for which the computed value $\Gamma(u_{\textnormal{test}},v)$ is positive.
Let $s=\arg\min_{\begin{subarray}{c}v\neq u_{\text{test}}\end{subarray}}\{\Gamma(u_{\text{test}},v)>0\}$, which corresponds to the user achieving the minimum positive value and hence defines $\hat{\gamma}$. Then, for any other user $v$, we must have either $\Gamma(u_{\text{test}},v)\leq 0$ or $\Gamma(u_{\text{test}},v)\geq\hat{\gamma}$.
We will show that any user included in $\mathcal{W}_{\hat{\gamma}}(u_{\text{test}})$, as defined in Theorem 4.2, also satisfies the condition in Lemma 4.7. For simplicity, we denote $u=u_{\text{test}}$ in the rest of the proof.
We first consider the case where $\Gamma(u,v)\leq 0$. This implies:
$\left\|\hat{\bm{\theta}}_{u}-\hat{\bm{\theta}}_{v}\right\|_{2}-\alpha(\text{CI}_{u}+\text{CI}_{v})\leq 0.$ (15)
By the triangle inequality, we have:
$\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}$ $\leq\left\|\hat{\bm{\theta}}_{u}-\hat{\bm{\theta}}_{v}\right\|_{2}+\|\hat{\bm{\theta}}_{u}-\bm{\theta}_{u}\|_{2}+\|\hat{\bm{\theta}}_{v}-\bm{\theta}_{v}\|_{2}$
$\leq\left\|\hat{\bm{\theta}}_{u}-\hat{\bm{\theta}}_{v}\right\|_{2}+\text{CI}_{u}+\text{CI}_{v}$
$\leq(\alpha+1)(\text{CI}_{u}+\text{CI}_{v}).$
Furthermore, using the definition of $K_{u}$ and $K_{v}$, we obtain:
$\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}\leq 2(\alpha+1)\sqrt{\max\{2,d\}}(K_{u}+K_{v}).$
This implies that
$K_{u}+K_{v}\geq\frac{\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}}{2(\alpha+1)\sqrt{\max\{2,d\}}}.$
Thus, by letting $\alpha_{r}=\frac{1}{2(\alpha+1)\sqrt{\max\{2,d\}}}$, we obtain the desired result:
$K_{u}+K_{v}\geq\alpha_{r}\|\bm{\theta}_{u}-\bm{\theta}_{v}\|_{2}.$
A.3 Proof of Lemma 4.8
Proof of Lemma 4.8.
The proof of Lemma 4.8 follows from the fact that for all $v\in M(u_{\textnormal{test}})$, we have
$\tilde{\Gamma}(u_{\textnormal{test}},v)\geq\|\bm{\theta}_{u_{\textnormal{test}}}-\bm{\theta}_{v}\|_{2}\geq\gamma.$
Therefore, as long as $M(u_{\textnormal{test}})\neq\emptyset$, the optimistic estimation ensures that $\hat{\gamma}$ is at least $\gamma$. ∎
A.4 Proof of Theorem 5.2
Proof of Theorem 5.2.
Let $\beta(n,\delta)=\sqrt{d\log\left(1+\frac{n}{\lambda d}\right)+2\log\left(\frac{1}{\delta}\right)}+\sqrt{\lambda}$, and let $\tilde{j}_{u}$ denote the cluster index of user $u$ in the clustering output $\tilde{\mathcal{G}}$. We denote by $\tilde{\mathcal{V}}(j)$ the $j$-th cluster in $\tilde{\mathcal{G}}$.
We first introduce the following events:
[…]
$\mathcal{F}_{u}(\delta)$ =$=\left\{\left\|\bm{\theta}_{u}-\tilde{\bm{\theta}}_{u}\right\|_{\tilde{M}_{u}}\leq\beta(\tilde{N}_{u},\delta)\right\},\quad\forall u\in\mathcal{V},$
$\mathcal{F}^{\prime}_{u}(\delta)$ =$=\left\{\left\|\bm{\theta}_{u}-\hat{\bm{\theta}}_{u}\right\|_{M_{u}}\leq\beta(N_{u},\delta)\right\},\quad\forall u\in\mathcal{V}.$ (16)
We first verify that event $\mathcal{E}$ holds with high probability. According to Lemma B.1, the event $\mathcal{F}^{\prime}_{u}(\delta/(2U))$ holds for all $u\in\mathcal{V}$ with probability at least $1-\frac{1}{2}\delta$. Under this event, we have:
$\left\|\hat{\bm{\theta}}_{u}-\bm{\theta}_{u}\right\|_{2}$ $\leq\frac{\left\|\hat{\bm{\theta}}_{u}-\bm{\theta}_{u}\right\|_{M_{u}}}{\sqrt{\lambda_{\min}(M_{u})}}$
$\leq\frac{\beta(N_{u},\delta/(2U))}{\sqrt{\lambda+\lambda_{\min}\left(\sum_{i=1}^{N_{u}}\bm{a}_{u}^{i}(\bm{a}_{u}^{i})^{\top}\right)}}$
$\leq\frac{\sqrt{d\log\left(1+\frac{N_{u}}{\lambda d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{\lambda}}{\sqrt{\lambda+\lambda_{a}N_{u}/2}}$
$\leq\alpha\cdot\mathrm{CI}_{u}\leq\frac{\gamma}{4},$ (17)
with probability at least $1-\frac{3}{2}\delta$, for any $u$ such that
$N_{u}\geq\max\left\{\frac{512d}{\gamma^{2}\lambda_{a}}\log\left(\frac{2U}{\delta}\right),\ \frac{16}{\lambda_{a}^{2}}\log\left(\frac{8dU}{\lambda_{a}^{2}\delta}\right)\right\}.$
The last inequality in (17) follows from Lemma B.4.
Now consider the connection rule in Algorithm 2, specifically the separation condition in Equation 7. For any users $u_{1},u_{2}\in\mathcal{V}$, the algorithm distinguishes them into different clusters if
$\alpha(\mathrm{CI}_{u_{1}}+\mathrm{CI}_{u_{2}})<\left\|\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}\right\|_{2}.$
From the triangle inequality, we have:
$\left\|\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}\right\|_{2}$ $\leq\left\|\bm{\theta}_{u_{1}}-\bm{\theta}_{u_{2}}\right\|_{2}+\left\|\hat{\bm{\theta}}_{u_{1}}-\bm{\theta}_{u_{1}}\right\|_{2}+\left\|\hat{\bm{\theta}}_{u_{2}}-\bm{\theta}_{u_{2}}\right\|_{2}$
$\leq\left\|\bm{\theta}_{u_{1}}-\bm{\theta}_{u_{2}}\right\|_{2}+\alpha\cdot\mathrm{CI}_{u_{1}}+\alpha\cdot\mathrm{CI}_{u_{2}}.$
Therefore, combining both bounds gives:
$\alpha(\mathrm{CI}_{u_{1}}+\mathrm{CI}_{u_{2}})<\left\|\bm{\theta}_{u_{1}}-\bm{\theta}_{u_{2}}\right\|_{2}+\alpha(\mathrm{CI}_{u_{1}}+\mathrm{CI}_{u_{2}}),$
which implies $\left\|\bm{\theta}_{u_{1}}-\bm{\theta}_{u_{2}}\right\|_{2}>0$, and since the true cluster gap is at least $\gamma$, we conclude:
$\left\|\bm{\theta}_{u_{1}}-\bm{\theta}_{u_{2}}\right\|_{2}\geq\gamma,$
i.e., $u_{1}$ and $u_{2}$ belong to different clusters. This confirms that the algorithm correctly separates users from different clusters with high probability, and hence event $\mathcal{E}$ holds with probability at least $1-\frac{3}{2}\delta$.
On the other hand, for users $u_{1},u_{2}$ that do not satisfy the separation condition, we have:
$\gamma$ $\geq 2\alpha(\mathrm{CI}_{u_{1}}+\mathrm{CI}_{u_{2}})\geq\left\|\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}\right\|_{2}+\alpha(\mathrm{CI}_{u_{1}}+\mathrm{CI}_{u_{2}})$
$\geq\left\|\bm{\theta}_{u_{1}}-\hat{\bm{\theta}}_{u_{1}}\right\|_{2}+\left\|\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}\right\|_{2}+\left\|\bm{\theta}_{u_{2}}-\hat{\bm{\theta}}_{u_{2}}\right\|_{2}\geq\left\|\bm{\theta}_{u_{1}}-\bm{\theta}_{u_{2}}\right\|_{2}.$
This implies that $u_{1}$ and $u_{2}$ must belong to the same cluster. Therefore, the clustering result satisfies event $\mathcal{E}$ with probability at least $1-\delta$.
Furthermore, due to the homogeneity of users within each cluster $\tilde{\mathcal{V}}(j)$, the event $\mathcal{F}^{\prime}_{u}(\delta/(2U))$ also holds for all $u\in\mathcal{V}$ with probability at least $1-\delta$.
We now analyze the suboptimality of Off-CLUB. Recall that:
$\text{SubOpt}(u_{\text{test}},\text{Off-CLUB},A_{\mathrm{test}})=\left\langle\bm{\theta}_{u},\bm{a}_{\text{test}}^{*}\right\rangle-\left\langle\bm{\theta}_{u},\bm{a}_{\text{test}}\right\rangle.$
When event $\mathcal{F}_{u}(\delta/(2U))$ holds for all $u\in\mathcal{V}$, by the Cauchy–Schwarz inequality we have:
$\left\langle\tilde{\bm{\theta}}_{u}-\bm{\theta}_{u},\bm{a}_{\text{test}}\right\rangle\leq\beta\left(\tilde{N}_{u},\delta/(2U)\right)\cdot\left\|\bm{a}_{\text{test}}\right\|_{\tilde{M}_{u}^{-1}}.$
Therefore,
$-\left\langle\bm{\theta}_{u},\bm{a}_{\text{test}}\right\rangle$ $\leq-\left\langle\tilde{\bm{\theta}}_{u},\bm{a}_{\text{test}}\right\rangle+\beta\left(\tilde{N}_{u},\delta/(2U)\right)\cdot\left\|\bm{a}_{\text{test}}\right\|_{\tilde{M}_{u}^{-1}}$
$\leq-\left\langle\tilde{\bm{\theta}}_{u},\bm{a}_{\text{test}}^{*}\right\rangle+\beta\left(\tilde{N}_{u},\delta/(2U)\right)\cdot\left\|\bm{a}_{\text{test}}^{*}\right\|_{\tilde{M}_{u}^{-1}}.$ (18)
Then, it follows that
$\text{SubOpt}(u,\pi,A_{\text{test}})$ $\leq\left(\bm{\theta}_{u}-\tilde{\bm{\theta}}_{u}\right)^{\top}\bm{a}_{\text{test}}^{*}+\beta\left(\tilde{N}_{u},\delta/(2U)\right)\cdot\left\|\bm{a}_{\text{test}}^{*}\right\|_{\tilde{M}_{u}^{-1}}$
$\leq\left\|\bm{a}_{\text{test}}^{*}\right\|_{\tilde{M}_{u}^{-1}}\cdot\left\|\tilde{\bm{\theta}}_{u}-\bm{\theta}_{u}\right\|_{\tilde{M}_{u}}+\beta\left(\tilde{N}_{u},\delta/(2U)\right)\cdot\left\|\bm{a}_{\text{test}}^{*}\right\|_{\tilde{M}_{u}^{-1}}$
$\leq 2\cdot\beta\left(\tilde{N}_{u},\delta/(2U)\right)\cdot\left\|\bm{a}_{\text{test}}^{*}\right\|_{\tilde{M}_{u}^{-1}},$
where the second inequality uses Cauchy–Schwarz, and the last follows from the definition of event $\mathcal{F}_{u}(\delta/(2U))$.
Note that
$\left\|\bm{a}_{\text{test}}^{*}\right\|_{\tilde{M}_{u}^{-1}}=\sqrt{\left(\bm{a}_{\text{test}}^{*}\right)^{\top}\tilde{M}_{u}^{-1}\bm{a}_{\text{test}}^{*}}=\sqrt{\left\langle\tilde{M}_{u}^{-1},\bm{a}_{\text{test}}^{*}\left(\bm{a}_{\text{test}}^{*}\right)^{\top}\right\rangle},$
and by definition $\tilde{M}_{u}=\lambda I+\sum_{(\bm{a},r)\in\tilde{\mathcal{D}}_{u}}\bm{a}\bm{a}^{\top}$.
From Lemma B.2, we know
$\lambda_{\min}(\tilde{M}_{u})\geq\lambda+\frac{\lambda_{a}N^{\mathcal{V}(j_{u})}}{2},$
and thus
$\left\|\bm{a}_{\text{test}}^{*}\right\|_{\tilde{M}_{u}^{-1}}\leq\frac{1}{\sqrt{\lambda+\lambda_{a}N^{\mathcal{V}(j_{u})}/2}}.$
Combining this with the definition of $\beta$, we conclude:
$\text{SubOpt}(u_{\text{test}},\text{Off-CLUB},A_{\text{test}})\leq\frac{\sqrt{d\log\left(1+\frac{N^{\mathcal{V}(j_{u_{\text{test}}})}}{\lambda d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{\lambda}}{\sqrt{\lambda+\lambda_{a}N^{\mathcal{V}(j_{u_{\text{test}}})}/2}},$ (19)
which holds with probability at least $1-2\delta$. Using the order $\tilde{O}$ to denote Equation 19 ends our proof. ∎
A.5 Proof of Theorem 6.1
Proof of Theorem 6.1.
Note that within the dataset $\mathcal{D}$, only the samples collected from users in the same cluster $\mathcal{V}(j_{u_{\text{test}}})$ are homogeneous with the test user. Therefore, the total number of homogeneous samples available for $u_{\text{test}}$ is
$N^{\mathcal{V}(j_{u_{\text{test}}})}=\sum_{v\in\mathcal{V}(j_{u_{\text{test}}})}N_{v}.$
By applying Theorem 2 from Li et al. [2022] with parameters $\Lambda=2\sqrt{2}$, $p=q=2$, and confidence bound term $\mathrm{CB}_{q}(\Lambda)=\mathcal{I}_{u_{\text{test}}}$, we directly obtain Theorem 6.1. This completes the proof.
Appendix B Technical Lemmas
We introduce several lemmas which are used in our proof.
Lemma B.1.
Suppose $(\bm{a}_{1},R),\dots,(\bm{a}_{n},r_{n}),\dots$ are generated sequentially from a linear model such that $\left\|\bm{a}_{s}\right\|_{2}\leq 1$ for all $s$, $\mathbb{E}[r_{s}|\bm{a}_{s}]=\bm{\theta}^{\top}\bm{a}_{s}$ for fixed but unknown $\bm{\theta}$, and $\{r_{s}-\theta^{\top}\bm{a}_{s}\}_{s=1,2,\cdots}$ have $R$-sub-Gaussian tails. Let $M_{n}=\lambda I+\sum_{s=1}^{n}\bm{a}_{s}\bm{a}_{s}^{\top}$, $\bm{b}_{n}=\sum_{s=1}^{n}r_{s}\bm{a}_{s}$ and $\delta>0$. Let $\hat{\bm{\theta}}_{s}=M_{s}^{-1}\bm{b}_{s}$ is the ridge regression estimator of $\bm{\theta}$. Then
$\left\|\hat{\bm{\theta}}_{n}-\bm{\theta}\right\|_{M_{n}}\leq R\sqrt{d\log\left(1+\frac{n}{\lambda d}\right)+2\log\left(\frac{1}{\delta}\right)}+\sqrt{\lambda}$
holds with probability at least $1-\delta$.
Lemma B.2.
Let $\bm{a}_{s},n\geq 1$ be generated sequentially from a random distribution such that $||\bm{a}||_{2}\leq 1$ and $\mathbb{E}[\bm{a}\bm{a}^{\top}]$ is full rank with minimal eigenvalue $\lambda_{a}>0$. Let $M_{n}=\sum_{s=1}^{n}\bm{a}_{s}\bm{a}_{s}^{\top}$. Then event
$\lambda_{\min}(M_{n})\geq\left(n\lambda_{a}-\frac{1}{3}\sqrt{18nA(\delta)+A(\delta)^{2}}-\frac{1}{3}A(\delta)\right)$
holds with probability at least $1-\delta$ for $n\geq 0$ where $A(n,\delta)=\log\left(\frac{(n+1)(n+3)d}{\delta}\right)$. Furthermore,
$\lambda_{\min}(M_{n})\geq\frac{1}{2}\lambda_{a}n,\ \forall n\geq\frac{16}{\lambda_{a}^{2}}\log\left(\frac{8d}{\lambda_{a}^{2}\delta}\right)$
holds with probability at least $1-\delta$.
Lemma B.3.
If $a,b>0$ and $ab>e$, then $\forall t\geq 2a\log(ab)$, we have $t\geq a\log(bt)$ holds.
Lemma B.4.
When $n\geq\frac{512d}{\gamma^{2}\lambda_{a}}\log(\frac{2U}{\delta})$ and $\lambda\leq d\log(1+\frac{n}{\lambda d})+2\log\left(\frac{2U}{\delta}\right)$,
$\frac{\sqrt{d\log(1+\frac{n}{\lambda d})+2\log(\frac{2U}{\delta})}+\sqrt{\lambda}}{\sqrt{\lambda_{a}\lambda+\lambda_{a}n/2}}\leq\frac{\gamma}{4}.$
The proofs of the above lemmas can be found in Li and Zhang [2018] and Wang et al. [2025]. Specifically, Lemma B.1 corresponds to Lemma 1, Lemma B.3 to Lemma 9, and Lemma B.4 to Lemma 10 in Li and Zhang [2018]. Lemma B.2 corresponds to Lemma 7 in Li and Zhang [2018] and Lemma B.2 in Wang et al. [2025].
Appendix C Additional Experiments
C.1 Basic Setup and Dataset Description
[…]
Dataset. (1). Synthetic Dataset: The user preference vectors $\bm{\theta}_{u}$ and item feature vectors $\bm{a}_{u}^{i}$ are in $d=20$-dimensional space, with each entry drawn from a standard Gaussian distribution and normalized to satisfy $\|\bm{\theta}_{u}\|_{2}=1$. (2). Real-world Datasets: We use two real-world recommendation datasets: We extract $1k$ users and $1k$ items with the highest interaction frequencies, then apply Singular Value Decomposition (SVD) to extract user preference vectors. For each user and each $i$, the action set $\mathcal{A}_{i}$ consists of 20 items, where each item’s feature vector is independently sampled from a sub-Gaussian distribution and then normalized to unit norm, following Li et al. [2019].
C.2 Benchmark Methods and Their Descriptions
To ensure fair comparisons, we adapt several online learning algorithms to the offline setting, modifying them as needed to work with offline datasets. Additionally, we propose an enhanced version of DBSCAN Schubert et al. [2017], tailored specifically to our environment. Compared to KMeans, DBSCAN is more suitable for our framework due to its parameterization (neighborhood radius and minimum samples), which aligns better with our requirements. For learning methods, We evaluate the following benchmarks:
  • SCLUB: A method that employs a hierarchical clustering structure to model users and enhances learning efficiency by optimizing cluster centers.
  • LinUCB: A linear regression-based Upper Confidence Bound (UCB) approach that dynamically balances exploration and exploitation for model selection.
  • LinUCB_Ind: An independent variant of LinUCB that operates without leveraging any clustering information. This method applies the LinUCB approach independently to each user.
  • CLUB: A cluster-based multi-armed bandit method that groups users from a complete graph according to their characteristics to improve model selection.
  • DBSCAN_Improve: Similar to CLUB, but uses DBSCAN for the clustering stage.
  • XMeans_Improve: Similar to CLUB, but uses XMeans for the clustering stage.
Figure 4: Suboptimality Gap under Varying $\hat{\gamma}$ with $|\mathcal{D}|=30000$.
Figure 5: Suboptimality Gap under Varying $\hat{\gamma}$ with $|\mathcal{D}|=35000$.
Figure 6: Suboptimality Gap under Varying $\hat{\gamma}$ with $|\mathcal{D}|=40000$.
Figure 7: Suboptimality Gap under Different User Probability Distributions and LinUCB Sampling Strategy with Insufficient Data.
C.3 The Dimension of Preference Vector
We first select $1k$ users and $1k$ items with the highest interaction frequencies, construct the binary feedback matrix based on users’ ratings. Then we use SVD to decompose the extracted binary feedback matrix $R_{u\times m}=\theta SX^{T}$, where $\theta=(\theta_{i}),i\in[u]$ and $X=(x_{j}),j\in[m]$. Note that the user $\bm{\theta}_{u}$ distribution in this setting does not fully match the assumptions from the simulation, as the user $\bm{\theta}_{u}$ values are scattered and exhibit clustering behavior.
We apply the Elbow Method to select the optimal dimension $d$. The objective is to identify the dimension that captures the most significant variance while ensuring computational efficiency. Figures 9 and 9 show the normalized variance explained by the first $d$ singular values for both datasets. In both cases, the "elbow" occurs at about $d=20$, where the marginal gain in explained variance diminishes. Thus, $d=20$ is selected as the optimal dimension for both datasets, achieving a balance between model complexity and variance preservation.
Figure 8: SVD Elbow Method for Yelp Dataset.
Figure 9: SVD Elbow Method for MovieLens Dataset.
C.4 User Probability Distribution (Detailed Analysis)
To determine the dataset size for each user, we consider two sampling strategies. When determining the scale of each $\mathcal{D}_{i}$ in offline data sampling, we mainly consider two distributions:
  • Equal Distribution: The probability of sampling data is uniform across all users, which gives
    $E(|\mathcal{D}_{i}|)=\frac{|\mathcal{D}|}{U}.$
  • Semi-Random Distribution: While the probability within each cluster is uniform, the probabilities of selecting different clusters vary. Each class has its own probability while users within the same class share the same probability. This can be expressed as
    $E(|\mathcal{D}_{i}|)=\frac{|\mathcal{D}|}{U}\cdot p_{J(i)},\quad p=[p_{1},p_{2},\ldots,p_{J}],\quad\sum_{i=1}^{J}p_{i}=1,\quad p_{i}\geq 0.$
Table 5: Impact of Increased Data on Suboptimality gap under Yelp Dataset.
Algorithm Dataset size $0.2M$ $0.4M$ $0.6M$ $0.8M$ $1M$
DBSCAN_Improve 0.146196 0.069701 0.033431 0.018931 0.011499
XMeans_improve 0.187885 0.138490 0.085507 0.098608 0.142219
CLUB 0.133169 0.037698 0.021715 0.013878 0.010067
Off-CLUB 0.022586 0.021187 0.012889 0.007229 0.004594
LinUCB 0.152432 0.151958 0.152372 0.151933 0.152491
LinUCB_IND 0.143812 0.066350 0.027777 0.013595 0.007608
SCLUB 0.176953 0.121705 0.041028 0.040673 0.041705
Off-C2LUB_Pessimistic 0.021467 0.018053 0.011366 0.007157 0.004992
Off-C2LUB_Optimistic 0.023665 0.020053 0.011782 0.007168 0.004709
Table 6: Impact of Increased Data on Suboptimality gap under MovieLens Dataset.
Algorithm Dataset size $0.2M$ $0.4M$ $0.6M$ $0.8M$ $1M$
OffDBSCAN_Improve 0.149179 0.072216 0.033121 0.015387 0.010090
OffXMeans_improve 0.200525 0.189838 0.178002 0.167926 0.159602
CLUB 0.147572 0.060904 0.026974 0.013053 0.007262
Off-CLUB 0.087765 0.042926 0.019609 0.010277 0.005976
LinUCB 0.159126 0.159954 0.159370 0.159623 0.159583
LinUCB_IND 0.145191 0.066178 0.028085 0.013468 0.007575
SCLUB 0.174606 0.117944 0.039357 0.039316 0.039230
Off-C2LUB_Pessimistic 0.083684 0.040423 0.019339 0.010653 0.006760
Off-C2LUB_Optimistic 0.089122 0.041868 0.019698 0.010243 0.005970
C.5 Offline Data Distribution (Detailed Analysis)
The offline dataset is influenced by the user’s strategy during the sampling phase. For each data point in the dataset $\mathcal{D}_{u}$, a subset of 20 items ($\mathcal{A}_{i}$) is generated for sampling. We consider two selection strategies:
  • Random Selection: The user randomly selects one item from $\mathcal{A}_{i}$, denoted as $a_{u}^{i}\sim\mathrm{Uniform}(\mathcal{A}_{i})$.
  • LinUCB Selection: During the offline phase, each user independently applies the LinUCB method Abbasi-Yadkori et al. [2011] to select one item from $\mathcal{A}_{i}$, denoted as $a_{u}^{i}=\mathop{\mathrm{argmax}}\limits_{a\in\mathcal{A}_{i}}\langle\theta_{u},x_{a}\rangle+\alpha\sqrt{x_{a}^{\top}M_{u}^{-1}x_{a}}$.
When data is collected randomly, it implies that users are selecting items without exhibiting explicit preferences. Under such circumstances, the collected data spans all feature dimensions uniformly, ensuring broad coverage but lacking targeted exploration.
By contrast, a more informed data collection strategy, such as LinUCB, allows the system to identify and concentrate on more informative dimensions. This offline process still relies on item-reward pairs from various arms, enabling the feature space to be explored in a more focused manner. Consequently, this approach yields tighter confidence intervals and enhances the estimation accuracy of the algorithm.
In summary, while random selection ensures broad coverage of the feature space, LinUCB provides a more targeted exploration, leading to improved algorithmic performance. From a practical standpoint, if an existing recommendation system has been running under LinUCB, it can directly transition to the Off-C2LUB algorithm while leveraging the previously collected data. On the other hand, if newly designed algorithms demand uniform data distributions, it may be necessary to re-run the system in order to collect data adhering to these uniform assumptions.
C.6 Performance Analysis
Figure 7 illustrates the evolution of the average suboptimality gap under different real-world dataset scenarios (Yelp and MovieLens) over varying dataset sizes ($|\mathcal{D}|$). For Off-C2LUB, $\hat{\gamma}$ is optimized in an equal distribution scenario and applied to semi-random distribution. The results indicate that, despite no changes in the distribution of users’ $\bm{\theta}_{u}$, The optimized $\hat{\gamma}$ demonstrates environment sensitivity, necessitating adaptive strategies for robust performance. This underscores the importance of the optimistic-pessimistic strategy, which dynamically adapts to varying environments to achieve better performance. In real-world datasets, the distribution probability of users is their inherent attribute. In this case, the optimal gamma obtained through search can still achieve good performance under different sampling strategies. Notably, in the severely insufficient data regime ($|\mathcal{D}|\in[20k,100k]$), Off-C2LUB_Optimistic(Pessimistic) achieves significant improvements over baseline methods.
Table C.4 and Table C.4 present experimental suboptimality gap demonstrating that our proposed methods achieve excellent convergence, even in scenarios with relatively abundant data. Compared to traditional approaches such as CLUB and SCLUB, our methods exhibit faster convergence and maintain a significantly smaller suboptimality gap, highlighting their efficiency and robustness. This reduction in the suboptimality gap underscores the superior performance of our methods in achieving near-optimal solutions under varying data conditions.
Appendix D Further Discussions on Off-CLUB
D.1 Detailed Algorithm: Off-CLUB
The pseudo-code of Off-C2LUB is as follows:
Algorithm 2 Off-CLUB
1:   Input: User $u_{\text{test}}$, action set $A_{\text{test}}$, dataset $\mathcal{D}$, parameters $\alpha$, $\lambda>1$ and $\delta>0$
2:   Initialization: Construct a complete graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, and compute $M_{u}$, $\bm{b}_{u}$, $\hat{\bm{\theta}}_{u}$, and $\text{CI}_{u}$ as in (1) for each user $u\in\mathcal{V}$
[…]
4:  for each edge $(u_{1},u_{2})\in\mathcal{E}$ do
5:      Remove edge $(u_{1},u_{2})$ if condition (7) is satisfied
6:  end for
7:  Let $\tilde{\mathcal{G}}=(\mathcal{V},\tilde{\mathcal{E}})$ denote the updated graph
8:  for each user $u\in\mathcal{V}$ do
9:      Aggregate data:
$\tilde{\mathcal{D}}_{u}=$ $\bigcup_{(u,v)\in\tilde{\mathcal{E}}}\left\{(\bm{a}_{v}^{i},r_{v}^{i})\in\mathcal{D}_{v}\right\}\cup\mathcal{D}_{u},$
10:      Compute cluster statistics:
$\tilde{M}_{u}=\lambda I+\sum_{(\bm{a},r)\in\tilde{\mathcal{D}}_{u}}\bm{a}\bm{a}^{\top},\quad\tilde{N}_{u}=\left|\tilde{\mathcal{D}}_{u}\right|,\quad\tilde{\bm{b}}_{u}=\sum_{(\bm{a},r)\in\tilde{\mathcal{D}}_{u}}r\bm{a},\quad\tilde{\bm{\theta}}_{u}=\left(\tilde{M}_{u}\right)^{-1}\tilde{\bm{b}}_{u}$
11:  end for
[…]
13:   Select the action:
$\bm{a}_{\text{test}}=\mathop{\mathrm{argmax}}\limits_{\bm{a}\in A_{\text{test}}}\left(\left(\tilde{\bm{\theta}}_{u_{\text{test}}}\right)^{\top}\bm{a}-\beta\left\|\bm{a}\right\|_{\left(\tilde{M}_{u_{\text{test}}}\right)^{-1}}\right)$
where $\beta=\sqrt{d\log\left(1+\frac{\tilde{N}_{u_{\text{test}}}}{\lambda d}\right)+2\log\left(\frac{2U}{\delta}\right)}+\sqrt{\lambda}$
The algorithm takes similar inputs to Algorithm 1, except that it does not require $\hat{\gamma}$. It starts with a complete graph $\mathcal{G}$ and iterates through all edges $(u_{1},u_{2})\in\mathcal{E}$ during the Cluster Phase. If the condition $\left\|\hat{\bm{\theta}}_{u_{1}}-\hat{\bm{\theta}}_{u_{2}}\right\|_{2}>\alpha\left(\text{CI}_{u_{1}}+\text{CI}_{u_{2}}\right)$ holds, the algorithm concludes that users $u_{1}$ and $u_{2}$ do not share the same preference vector and removes the edge between them (Line 5). After processing all edges, the resulting graph $\tilde{\mathcal{G}}$ separates users with different preference vectors into distinct clusters, provided that each user has sufficient samples to ensure a low confidence interval. However, if a user $u$ has fewer samples in $\mathcal{D}_{u}$, the resulting higher confidence interval $\text{CI}_{u}$ reduces the likelihood of satisfying condition (7), potentially leading to the inclusion of heterogeneous users in the same cluster. The conditions necessary to ensure the algorithm’s correctness are discussed in detail in Section 5.2.
Similar to Algorithm 1, this algorithm aggregates data from each user and its neighbors in $\tilde{\mathcal{G}}$ to form a new dataset and computes the necessary statistics (Lines 9 and 10). It is important to note that this data aggregation strategy, like that in Algorithm 1, differs from the online version of the CLUB algorithm introduced in [Gentile et al., 2014], where data from all users within the same connected component of the graph is utilized. For instance, consider a scenario where $(u,v)\in\tilde{\mathcal{E}}$ and $(v,w)\in\tilde{\mathcal{E}}$, but $(u,w)\notin\tilde{\mathcal{E}}$. In this case, samples from $\mathcal{D}_{w}$ are excluded from $\tilde{\mathcal{D}}_{u}$ in Algorithm 2, whereas the online version includes them in $\tilde{\mathcal{D}}_{u}$. The restricted aggregation in Algorithm 2 is crucial in the offline setting: since users that are no longer neighbors of $u$ are guaranteed not to share the same preference vector, including their data in $\tilde{\mathcal{D}}_{u}$ would introduce bias. In the online setting, however, such bias is not a concern because infinite online data ensures that all edges between users with different preference vectors are reliably removed. This distinction highlights the importance of careful data selection in the offline setting to maintain prediction accuracy.
Finally, the Decision Phase is performed in a manner analogous to Algorithm 2, using a pessimistic estimate to select $\bm{a}_{\text{test}}$.
D.2 Additional Remarks
Remark D.1 (Asymptotic Properties).
In the special case where $\mathcal{V}(j_{u_{\textnormal{test}}})=\{u_{\textnormal{test}}\}$ (i.e., the test user is distinct from all other users in $\mathcal{V}$), the result reduces to the standard bound for linear bandits, $\tilde{O}\left(\sqrt{\frac{d}{N_{u_{\textnormal{test}}}}}\right)$. Conversely, when $\mathcal{V}(j_{u_{\textnormal{test}}})\approx\mathcal{V}$ (i.e., most users share the same preference vector), Algorithm 2 effectively aggregates data from nearly all users, leveraging the large sample size to make more accurate predictions.
Remark D.2 (Intuition of Assumption 5.1).
The key assumption underlying Theorem 5.2 is data sufficiency. As defined in Assumption 5.1, complete data sufficiency comprises two main components. The first component depends on $1/\lambda_{a}^{2}$, which ensures sufficient information about each dimension of the preference vector for each user. Without this level of data, the algorithm cannot reliably estimate each dimension, making accurate predictions infeasible. In most cases, $\lambda_{a}$ can be treated as a constant, rendering the first term effectively constant. The second component in Assumption 5.1 depends on $1/\gamma^{2}$, which is the main part of the data sufficiency requirement. This term guarantees that the true cluster of each user can be accurately identified. Insufficient data in this regard prevents the algorithm from correctly assigning users to their respective clusters, potentially leading to the inclusion of users with different preference vectors and introducing bias into the predictions. Notably, a smaller $\gamma$ corresponds to a stricter data sufficiency requirement.